### v1

# Towards an Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning

Abstract
As a marriage between offline RL and meta-RL, the advent of offline meta-reinforcement learning (OMRL) has shown great promise in enabling RL agents to multi-task and quickly adapt while acquiring knowledge safely. Among which, Context-based OMRL (COMRL) as a popular paradigm, aims to learn a universal policy conditioned on effective task representations. In this work, by examining several key milestones in the field of COMRL, we propose to integrate these seemingly independent methodologies into a unified information theoretic framework. Most importantly, we show that the pre-existing COMRL algorithms are essentially optimizing the same mutual information objective between the task variable $\bm{M}$ and its latent representation $\bm{Z}$ by implementing various approximate bounds. Based on the theoretical insight and the information bottleneck principle, we arrive at a novel algorithm dubbed UNICORN, which exhibits remarkable generalization across a broad spectrum of RL benchmarks, context shift scenarios, data qualities and deep learning architectures, attaining the new state-of-the-art. We believe that our framework could open up avenues for new optimality bounds and COMRL algorithms.
1 Introduction
The ability to swiftly learn and generalize to new tasks is a hallmark of human intelligence. In pursuit of this high level of artificial intelligence (AI), the paradigm of meta-reinforcement learning (RL) proposes to train AI agents in a trial-and-error manner by interacting with multiple external environments. In order to quickly adapt to the unknown, the agents need to integrate prior knowledge with minimal experience (namely the context) collected from the new tasks or environments, without over-fitting to the new data. This meta-RL mechanism has been adopted in many applications such as games (Zhao et al., 2022a; Beck et al., 2023), robotics (Yu et al., 2020a; Kumar et al., 2021) and drug discovery (Feng et al., 2022).
However, for data collection, classical meta-RL usually requires enormous online explorations of the environments (Finn et al., 2017; Rakelly et al., 2019), which is impractical in many safety-critical scenarios such as healthcare (Gottesman et al., 2019) and autonomous driving (Shalev-Shwartz et al., 2016). As a remedy, offline RL (Levine et al., 2020) enables agents to learn from logged experience only, therefore circumventing risky or costly online interactions. Recently, offline meta-RL (OMRL) (Li et al., 2021b; Mitchell et al., 2021; Dorfman et al., 2021) has emerged as a novel paradigm to significantly extend the applicable range of RL by "killing two birds in one stone": it builds powerful agents that can quickly learn and adapt by meta-learning, while leveraging offline RL mechanism to ensure a safe and efficient optimization procedure. In the context of supervised or self-supervised learning, which is de facto offline, OMRL is reminiscent of the multi-task learning (Raffel et al., 2020), meta-training (Finn et al., 2017) and fine-tuning (Hu et al., 2021; Li and Liang, 2021; Lester et al., 2021) of pre-trained large models. We envision it to be a cornerstone of RL foundation models (Bommasani et al., 2021; Reed et al., 2022) in the future.
Along the line of OMRL research, context-based offline meta-reinforcement learning (COMRL) is a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs). Intuitively, the crux of COMRL is learning effective task representations, and therefore enabling the agent to react optimally and adaptively in various contexts. To achieve this goal, one of the earliest COMRL algorithms FOCAL (Li et al., 2021b) proposes to capture the structure of task representations by distance metric learning. From a geometric perspective, it essentially performs clustering by pushing away latent embeddings of different tasks while pulling together those from the same task, hence ensuring consistent and distinguishable task representations.
Despite its effectiveness, FOCAL is reported to be vulnerable to context shifts (Li et al., 2021a), i.e., when testing on out-of-distribution (OOD) data. Such problems are particularly challenging for OMRL, since any context shift incurred at test time can not be rectified in the fully offline setting, which may result in severely degraded generalization performance (Li et al., 2021a; Gao et al., 2023). To alleviate the problem, follow-up works such as CORRO (Yuan and Lu, 2022) reformulates the task representation learning of COMRL as maximizing the mutual information $I(\bm{Z};\bm{M})$ between the task variable $\bm{M}$ and its latent $\bm{Z}$. It then approximates $I(\bm{Z};\bm{M})$ by an InfoNCE (Oord et al., 2018) contrastive loss, where the positive and negative pairs are conditioned on identical state-action tuples $(\bm{s},\bm{a})$. Inspired by CORRO, a recently proposed method CSRO (Gao et al., 2023) introduces an additional mutual information term between $\bm{Z}$ and $(\bm{s},\bm{a})$. By explicitly minimizing it along with the FOCAL objective, CSRO is demonstrated to achieve the state-of-the-art (SOTA) generalization performance on various MuJoCo (Todorov et al., 2012) benchmarks.
In this paper, following the recent development and storyline of COMRL, we present a novel and unified view of several signature methods including FOCAL, CORRO and CSRO under a information theoretic framework. We first show that these algorithms are essentially optimizing different approximate bounds of the same mutual information objective $I(\bm{Z},\bm{M})$. Then based on our proposed Unified Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning (UNICORN), we arrive at a new COMRL optimization scheme by deriving a unified and generalized task representation learning objective in light of the information bottleneck principle  (Tishby et al., 1999; Tishby and Zaslavsky, 2015). We demonstrate experimentally that the proposed method achieves state-of-the-art in-distribution and OOD generalization performance on a wide range of RL domains, OOD settings, data qualities and model architectures. Our framework potentially provides a principled roadmap to new COMRL algorithms by seeking better optimality bounds/approximations of the objective, as well as new implementations to further combat context shift.
2 Related Works
Offline Meta-Reinforcement Learning.   OMRL solves the problems of generalization and expensive data collection by meta-learning on pre-collected offline datasets. To overcome the overestimation of out-of-distribution actions (Kumar et al., 2019; Levine et al., 2020; Gao et al., 2022), offline RL usually leverages regularizations such as KL-divergence (Jaques et al., 2019), maximum mean discrepancy (MMD) (Kumar et al., 2019) and Wasserstein Distance (Gulrajani et al., 2017) to constrain the distribution mismatch between the behavior and learner policies. Moreover, a general mechanism called behavior regularized actor critic (BRAC) (Wu et al., 2019) involving value penalty and behavior regularization has proven successful in several previous studies (Li et al., 2021b; Gao et al., 2023), and is adopted in our method.
From a meta-learning perspective, the realization of OMRL can be categorized as gradient-based and context-based paradigms. Gradient-based meta-RL methods (Finn et al., 2017; Mitchell et al., 2021) embed an optimization procedure such as gradient descent into the meta-level iteration loops, which are model-agnostic but tend to reach lower asymptotic performance with less computational efficiency in practice. COMRL conditions the models on a contextual representation of MDP (Hallak et al., 2015) generated by recurrent (Wang et al., 2016; Fakoor et al., 2020), recursive (Mishra et al., 2017; Xu et al., 2022), probabilistic (Rakelly et al., 2019; Zintgraf et al., 2019) or deterministic (Li et al., 2021b, a; Yuan and Lu, 2022) task inference modules. At test time, it leverages neural networks as universal value/policy function approximators (Schaul et al., 2015) to perform meta-adaptation by transfer learning of the task representations. Depending on the source of the testing data, fully offline (Li et al., 2021b, a; Yuan and Lu, 2022) and offline-to-online (Zhao et al., 2022b; Pong et al., 2022) COMRL have been investigated. Notwithstanding, both paradigms may suffer significant OOD as the data-collecting policies might drastically differ from the training sets. In this work, we restrict our attention to fully offline COMRL, and demonstrate the remarkable robustness of our proposed framework to context shift in this setting.
3 Method
Figure 1: Context shift of COMRL in Ant-Dir. Left: Given a task specified by a goal direction (dashed line), the RL agent is trained on data generated by a variety of behavior policies trained on the same task (red). At test time, however, the context might be collected by behavior policies trained on different tasks (blue), causing a context shift of OOD behavior policies (Section 4.4). Middle: Under OOD testing context, UNICORN (red) is more robust than baselines such as FOCAL (green) in terms of navigating the Ant robot along the right direction. Right: Besides behavior policy, the task distribution (e.g., goal positions in Ant) can induce significant context shift (Section 5.2), which is also a challenging scenario for COMRL models to generalize.
3.1 Preliminaries and Problem Statement
We consider MDP modeled as $M=(\mathcal{S},\mathcal{A},T,R,\mathcal{\rho}_{0},\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $T(\bm{s}^{\prime}|\bm{s},\bm{a})$, bounded reward function $R(\bm{s},\bm{a})$, initial state distribution $\rho_{0}(\bm{s})$ and discount factor $\gamma\in(0,1)$. The goal is to find a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ to maximize the expected return. Starting from the initial state, for each time step, the agent performs an action sampled from $\pi$, then the environment updates the state with $T$ and returns a reward with $R$. We denote the marginal state distribution at time $t$ as $\mu_{\pi}^{t}(\bm{s})$. The V-function and Q-function are given by
$V_{\pi}(\bm{s})$ $=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\bm{s}_{t}\sim\mu_{\pi}^{t}(\bm{s}),\bm{a}_{t}\sim\pi}[R(\bm{s}_{t},\bm{a}_{t})]$ (1)
$Q_{\pi}(\bm{s},\bm{a})$ $=R(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s}^{\prime}\sim T(\bm{s}^{\prime}|\bm{s},\bm{a})}[V_{\pi}(\bm{s}^{\prime})].$ (2)
In this paper, we restrict to scenarios where tasks share the same state and action space while differing in reward and transition functions, with each OMRL task defined as an instance of MDP: $M^{i}=(\mathcal{S},\mathcal{A},T^{i},R^{i},\rho_{0},\gamma)\in\mathcal{M}$, where $\mathcal{M}$ is the set of all possible MDPs to be considered. For each task index $i$, the offline dataset $X^{i}=\{(\bm{s}_{j}^{i},\bm{a}_{j}^{i},r_{j}^{i},\bm{s}{{}^{\prime}}_{j}^{i})\}$ is collected by a behavior policy $\pi^{i}_{\beta}$. For meta-learning, given a trajectory segment $\bm{c}^{i}_{1:n}=\{(\bm{s}_{j}^{i},\bm{a}_{j}^{i},r_{j}^{i},\bm{s}{{}^{\prime}}_{j}^{i})\}_{j=1}^{n}$ of length $n$ as the context of $M^{i}$, COMRL employs a context encoder $q_{\bm{\phi}}(\bm{z}|\bm{c}_{1:n}^{i})$ to obtain the latent representation $\bm{z}^{i}$ of task $M^{i}$. If $\bm{z}$ contains enough information about the task identity, COMRL can be treated as a special case of RL on partially-observed MDP (Kaelbling et al., 1998), where $\bm{z}$ is interpreted as a faithful representation of the unobserved state. By conditioning on $\bm{z}$, the learning of universal policy $\pi_{\bm{\theta}}(\cdot|\bm{s},\bm{z})$ and value function $V_{\pi}(\bm{s},\bm{z})$ (Schaul et al., 2015) become regular RL, and optimality can be attained by Bellman updates with theoretical guarantees. To this end, FOCAL (Li et al., 2021b) proposes to decouple COMRL into the upstream task representation learning and downstream offline RL. For the former, FOCAL effectively clusters task embeddings by metric learning:
$\mathcal{L}_{\text{FOCAL}}=$ $\,\,\underset{\bm{\phi}}{\min}\,\,\mathbb{E}_{i,j}\text{\LARGE\{}\textbf{1}\{i=j\}||\bm{z}^{i}-\bm{z}^{j}||_{2}^{2}+\textbf{1}\{i\neq j\}\frac{\beta}{||\bm{z}^{i}-\bm{z}^{j}||_{2}^{n}+\epsilon}\text{\LARGE\}}.$ (3)
However, it is empirically shown that FOCAL struggles to generalize in the presence of context shift (Li et al., 2021a). This problem has been formulated in several earliest studies of OMRL.  Li et al. (2020) observed that when there are large divergence between the state-action distributions of the offline datasets, due to shortcut learning (Geirhos et al., 2020), the task encoder may learn to ignore the causal information like rewards and spuriously correlate primarily state-action pairs to the task identity, leading to poor generalization performance.  Dorfman et al. (2021) concurrently identified a related problem which they termed MDP ambiguity in the context of Bayesian offline RL. An example is illustrated in Figure 1. The problem is further exacerbated in the fully offline setting, as the testing distribution is fixed and cannot be augmented by online exploration, allowing no theoretical guarantee for the context shift.
To mitigate context shift, a subsequent algorithm CORRO (Yuan and Lu, 2022) proposes to optimize a lower bound of $I(\bm{Z};\bm{M})$ in the form of an InfoNCE (Oord et al., 2018) contrastive loss
$\mathcal{L}_{\text{CORRO}}$ $=\underset{\bm{\phi}}{\min}\,\,\mathbb{E}_{\bm{x},\bm{z}}\left[-\log\left(\frac{h(\bm{x},\bm{z})}{\sum_{M^{*}\in\mathcal{M}}h(\bm{x}^{*},\bm{z})}\right)\right],$ (4)
where $\bm{x}=(\bm{s},\bm{a},\bm{r},\bm{s}^{\prime})$, $\bm{z}\sim q_{\bm{\phi}}(\bm{z}|\bm{x})$, $h(\bm{x},\bm{z})=\frac{P(\bm{z}|\bm{x})}{P(\bm{z})}\approx\frac{q_{\bm{\phi}}(\bm{z}|\bm{x})}{P(\bm{z})}$, and $\bm{x}^{*}=(\bm{s},\bm{a},r^{*},\bm{s}^{\prime*})$ as a transition tuple generated for task $M^{*}\in\mathcal{M}$ conditioned on the same $(\bm{s},\bm{a})$. To further combat the spurious correlation between the task representation and behavior-policy-induced state-actions, a recently proposed method CSRO (Gao et al., 2023) introduces a CLUB upper bound (Cheng et al., 2020) of the mutual information between $\bm{z}$ and $(\bm{s},\bm{a})$, to regularize the FOCAL objective:
$\mathcal{L}_{\text{CSRO}}$ $=\underset{\bm{\phi}}{\min}\left\{\mathcal{L}_{\text{FOCAL}}+\lambda L_{\text{CLUB}}\right\}$ (5)
$L_{\text{CLUB}}$ $=\mathbb{E}_{i}\left[\log q_{\bm{\phi}}(\bm{z}_{i}|\bm{s}_{i},\bm{a}_{i})-\mathbb{E}_{j}\left[\log q_{\bm{\phi}}(\bm{z}_{j}|\bm{s}_{i},\bm{a}_{i})\right]\right].$ (6)
In the next section, we will show how these algorithms are inherently connected and can be further improved under our proposed unified information theoretic framework.
3.2 A Unified Information Theoretic Framework
We start with a formal definition of task representation learning in COMRL:
Definition 3.1.
Given an input context variable $\bm{X}\in\mathcal{X}$ and its associated task/MDP random variable $\bm{M}\in\mathcal{M}$, task representation learning in COMRL aims to find a minimal sufficient statistics $\bm{Z}$ of $\bm{X}$ with respect $\bm{M}$.
In pure statistical terms, Definition 3.1 implies that an ideal representation $\bm{Z}$ is the simplest mapping of $\bm{X}$ that captures the mutual information $I(\bm{X};\bm{M})$. We therefore assume the following dependency structures in terms of directed graphical models:
Assumption 3.2.
The dependency graphs of COMRL are given by
satisfying $I(\bm{Z};\bm{M};\bm{X}_{b})\geq 0$, where $\bm{X}_{b}$ and $\bm{X}_{t}$ are the behavior-related $(\bm{s},\bm{a})$-component and task-related $(\bm{s}^{\prime},r)$-component of the context $\bm{X}$, with $\bm{X}=(\bm{X}_{t},\bm{X}_{b})$.
For the first graph, $\bm{M}\rightarrow\bm{X}\rightarrow\bm{Z}$ forms a Markov chain, which satisfies $I(\bm{Z};\bm{M}|\bm{X})=0$. To interpret the second graph, given an MDP $M\sim\bm{M}$, the state-action component of $\bm{X}$ is fully captured by the behavior policy $\pi_{\beta}$: $\bm{s}\sim\mu_{\pi_{\beta}}(\bm{s}),\bm{a}\sim\pi_{\beta}$. We therefore define it as the behavior-related component, which should be minimally causally related (dashed lines) to $\bm{M}$ and $\bm{Z}$. Moreover, when $\bm{X}_{b}$ is given, $\bm{X}_{t}$ is fully characterized by the transition function $T:(\bm{s},\bm{a})\rightarrow\bm{s}^{\prime}$ and reward function $R:(\bm{s},\bm{a})\rightarrow r$ of $M$, which should be maximally causally related (solid lines) to $\bm{M}$ and $\bm{Z}$ and therefore be defined as the task-related component. Accordingly, we name $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ and $I(\bm{Z};\bm{X}_{b})$ the primary and lesser causality in our problem respectively. With this prior knowledge and assumption, we present the central theorem of this paper:
Theorem 3.3.
Let $\equiv$ denote equality up to a constant,
$\underbrace{I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})}_{\text{primary causality}}\quad\leq\quad I(\bm{Z};\bm{M})\quad\leq\underbrace{I(\bm{Z};\bm{X})}_{\text{primary + lesser causality}}$
holds up to a constant, where
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
[…]
Proof.
See Appendix B, Theorem B.1. ∎
Theorem 3.3 induces several key observations. Firstly, the FOCAL and CORRO objective operates as an upper bound and a lower bound of the relevant information $I(\bm{Z};\bm{M})$ respectively. Since one would like to maximize $I(\bm{Z};\bm{M})$ according to Definition 3.1, CORRO, which maximizes a lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$, can effectively optimize $I(\bm{Z};\bm{M})$ with theoretical guarantees. However, FOCAL which maximizes an upper bound $I(\bm{Z};\bm{X}_{t},\bm{X}_{b})$ is not guaranteed to optimize $I(\bm{Z};\bm{M})$. Mathematically since $I(\bm{Z};\bm{X})=I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+I(\bm{Z};\bm{X}_{b})$, maximizing the FOCAL objective may also increase the lesser causality $I(\bm{Z};\bm{X}_{b})$, which is not ideal since it contains spurious correlation between the task representation $\bm{Z}$ and behavior policy $\pi_{\beta}$. This explains why FOCAL is less robust to context shift compared to CORRO.
Secondly, CSRO as a latest COMRL algorithm, inherently optimizes a linear combination of the FOCAL and CORRO objectives. In the $0\leq\lambda\leq 1$ regime, the CSRO objective becomes a convex interpolation of the upper bound $I(\bm{Z};\bm{X})$ and the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ of $I(\bm{Z};\bm{M})$, which may form a tighter bound on the real objective $I(\bm{Z};\bm{M})$. This explains the improved performance of CSRO compared to FOCAL and CORRO.
3.3 UNICORN: Implementation of The Information Bottleneck Principle on COMRL
Given Definition 3.1 of the task representation learning, we now further generalize the pre-existing COMRL algorithms based on the information bottleneck principle (Tishby et al., 2000; Tishby and Zaslavsky, 2015). Namely, finding an optimal task representation $\bm{Z}$ can be formulated as the minimization of the following Lagrangian
$\mathcal{L}_{IB}=I(\bm{Z};\bm{X})-\beta I(\bm{Z};\bm{M})$ (7)
subject to the Markov chain constraint. The positive Lagrange multiplier $\beta$ acts as a tradeoff parameter between the complexity/rate of the representation, $R=I(\bm{Z};\bm{X})$, and the amount of preserved relevant information, $I_{M}=I(\bm{Z};\bm{M})$. For implementation, the first term is just FOCAL loss by Theorem 3.3, whereas we propose to approximate the second term as a convex combination of the FOCAL and CORRO loss, analogous to CSRO:
$I(\bm{Z};\bm{M})\approx\alpha I(\bm{Z};\bm{X})+(1-\alpha)I(\bm{Z};\bm{X}_{t}|\bm{X}_{b}),$ (8)
Figure 2: Meta-training procedure of UNICORN. Built upon the classic COMRL methods FOCAL, UNICORN implements the information bottleneck principle by jointly optimizing the FOCAL loss and a reconstruction loss, using a variational auto-encoder to obtain robust task representation $\bm{z}$.
where $0\leq\alpha\leq 1$. Substituting Eqn 8 into 7, we may rescale $\mathcal{L}_{IB}$ as
$\mathcal{L}_{IB}=-\left(I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+\frac{\alpha\beta-1}{(1-\alpha)\beta}I(\bm{Z};\bm{X})\right).$ (9)
For implementation of $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$, however, it would require additional approximations such as generating negative samples for contrastive learning using a trained model in CORRO, or using a CLUB upper bound of the mutual information in CSRO. Instead, we propose to rewrite it according to the chain rule of mutual information (Yeung, 2008):
$I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ $=I(\bm{X}_{t};\bm{Z},\bm{X}_{b})-I(\bm{X}_{t};\bm{X}_{b})$
$\equiv I(\bm{X}_{t};\bm{Z},\bm{X}_{b}),$ (10)
since $I(\bm{X}_{t};\bm{X}_{b})$ is a constant when $\bm{X}_{t}$ and $\bm{X}_{b}$ are drawn from a fixed distribution as in offline RL. Moreover, by definition of mutual information:
$I(\bm{X_{t}};\bm{Z},\bm{X_{b}})$ $=\int p(\bm{x_{t}},\bm{x_{b}})p(\bm{z}|\bm{x_{t}},\bm{x_{b}})\log\frac{p(\bm{x_{t}}|\bm{z},\bm{x_{b}})}{p(\bm{x_{t}})}$
$\equiv\int p(\bm{x_{t}},\bm{x_{b}})p(\bm{z}|\bm{x_{t}},\bm{x_{b}})\log p(\bm{x_{t}}|\bm{z},\bm{x_{b}})$
$\simeq\int_{\bm{x_{t}},\bm{x_{b}},\bm{z}}\underbrace{q_{\bm{\phi}}(\bm{z}|\bm{x_{t}},\bm{x_{b}})}_{\text{encoder}}\underbrace{\log p_{\bm{\theta}}(\bm{x_{t}}|\bm{z},\bm{x_{b}})}_{\text{decoder}}$
$=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x_{t}},\bm{x_{b}})}\left[\log p_{\bm{\theta}}(\bm{x_{t}}|\bm{z},\bm{x_{b}})\right],$ (11)
which induces an unbiased estimator of $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ by reconstructing $\bm{X}_{t}$ with a decoder network $p_{\bm{\theta}}(\cdot|\bm{z},\bm{x_{b}})$ conditioning on $\bm{Z}$ and $\bm{X_{b}}$, reminiscent of VAE (Kingma and Welling, 2014). We therefore implement the objective of our proposed algorithm UNICORN as a linear combination of this reconstruction loss $\mathcal{L}_{\text{recon}}:=-I(\bm{X_{t}};\bm{Z},\bm{X_{b}})$ and the FOCAL loss $\mathcal{L}_{\text{FOCAL}}=-I(\bm{Z};\bm{X})$:
$\mathcal{L}_{\text{UNICORN}}$ $:=\mathcal{L}_{IB}=\mathcal{L}_{\text{recon}}+\frac{\alpha\beta-1}{(1-\alpha)\beta}\mathcal{L}_{\text{FOCAL}}$
$=-\left(I(\bm{X}_{t};\bm{Z},\bm{X}_{b})+\frac{\alpha\beta-1}{(1-\alpha)\beta}I(\bm{Z};\bm{X})\right).$ (12)
UNICORN can be interpreted as a generalization of the FOCAL and CORRO objecitves by effectively striking a balance between the primary causality $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ and the lesser causality $I(\bm{Z};\bm{X}_{b})$. Figure 2 gives an overview of our framework, which enjoys three core advantages:
  • Robustness and bias-reduction. By optimizing a reconstruction loss $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ equivalent to the CORRO objective $I(\bm{Z};\bm{X_{t}}|\bm{X_{b}})$, UNICORN mitigates the spurious correlation in FOCAL, leading to more robust task representation $\bm{Z}$. It is also verified in Section 4.5 that UNICORN is robust to a wide range of data qualities. Moreover, the unbiased estimator in Eqn 11 circumvents additional approximations as in CORRO and CSRO, making UNICORN a bias-reduced and improved version of previous SOTAs.
  • Broader applicability domain. Unlike FOCAL/CORRO which assumes uniform task distribution, the unbiased estimator in Eqn 11 can be efficiently implemented by Monte Carlo sampling on offline datasets drawn from any distribution. Additionally, we will show in Section 5.1 that UNICORN is transferable to other model architectures such as Decision Transformer (DT).
  • General and unified paradigm. Derived from the general information bottleneck principle, our framework encompasses a series of milestones in COMRL including FOCAL, CORRO, CSRO and a concurrent method GENTLE (Zhou et al., 2023), which solely optimizes the reconstruction loss with a similar implementation. A unified picture of these pre-existing methods being special cases of UNICORN in the $\alpha$-$\beta$ plane is depicted in Figure 3. A hyper-parameter study showing the effect of the UNICORN feasible region is presented in Appendix C.1. Moreover, the decoder $p_{\theta}(\bm{x}_{t}|\bm{z},\bm{x}_{b})$ can operate as a model of the MDP to generate additional data using customized latent $\bm{z}$, which is explored in Section 5.2. In summary, UNICORN not only provides a unified framework of model-free COMRL, it may also be exploited and lead to novel model-based algorithms.
4 Experiments
Our main experiments are organized to address three primary inquiries regarding the core advantages of UNICORN: (1) How does UNICORN perform on in-distribution tasks (bias-reduction)? (2) How well can UNICORN generalize to data collected by OOD behavior policies (robustness)? (3) Can UNICORN outperform consistently across datasets of different qualities (robustness)? The performance are measured by the average return across all test tasks.
Figure 3: The hyper-parameter plane of UNICORN. Empirically, we find that UNICORN works properly only when $(\alpha\beta-1)/(1-\alpha)\beta>0$, which further constrains the feasible region of UNICORN (shaded) to be bounded by $\alpha\beta\geq 1$, $\alpha\leq 1$ and $\beta\geq 1$. Intriguingly, pre-existing methods FOCAL (red) and CORRO/CSRO/GENTLE (blue) constitute the two boundaries of the feasible region, which can therefore be interpreted as special cases of UNICORN.
4.1 Environments and Datasets
We adopt MuJoCo (Todorov et al., 2012) and MetaWorld (Yu et al., 2020a) benchmarks to evaluate our method.
HalfCheetah-Dir, HalfCheetah-Vel, Ant-Dir are multi-task MuJoCo benchmarks where tasks differ in reward function. For HalfCheetah-Dir and Ant-Dir, the agent is rewarded with high velocity along the goal direction. For HalfCheetah-Vel, the agent receives an $L_{2}$ penalty with respect to the target velocity.
Hopper-Param, Walker-Param are multi-task MuJoCo benchmarks where tasks differ in transition dynamics. The physical parameters of body mass, inertia, damping, and friction vary for each task.
Reach is a multi-task MetaWorld benchmark where tasks differ in reward function. The agent controls the end-effector to reach a target position in 3D space.
Following the protocol of CORRO, we sample 20 training tasks and 20 testing tasks from the task distribution. We train SAC agent (Haarnoja et al., 2018) independently for each task and collect offline datasets of mixed quality using all logged behavior policies.
4.2 Baselines
We compare UNICORN with six competitive OMRL algorithms. These baselines are categorized as context-based, gradient-based or transformer-based methods.
FOCAL, CORRO, CSRO are context-based methods which can be seen as special cases or approximations of UNICORN, see details in Section 3.
Supervised is a context-based method, directly using actor-critic loss to train the policy/value networks and task encoder end-to-end. We find it a competitive baseline across all benchmarks.
MACAW (Mitchell et al., 2021) is a gradient-based method, using supervised advantage-weighted regression for both the inner and outer loop of meta-training.
Prompt-DT (Xu et al., 2022) is a transformer-based method, taking context as the prompt of a Decision Transformer (DT) (Chen et al., 2021; Janner et al., 2021) to solve OMRL as a conditional sequence modeling problem.
We perform our evaluation in a few-shot manner. For gradient-based methods, we first update the meta-policy with a batch of context, then evaluate the agent. For transformer-based methods, we utilize $\bm{z}$ inferred from a batch of context as the prompt to condition the auto-regressive rollout.
4.3 Few-Shot Generalization to In-Distribution Data
Figure 4: Testing returns of UNICORN against baselines on six benchmarks. Solid curves refer to the mean performance of trials over 6 random seeds, and the shaded areas characterize the standard deviation of these trials.
Table 1: Average testing returns of UNICORN against baselines on datasets collected by IID and OOD behavior policies. Each result is averaged by 6 random seeds. The best is bolded and the second best is underlined.
Algorithm HalfCheetah-Dir HalfCheetah-Vel Ant-Dir Hopper-Param Walker-Param Reach
IID OOD IID OOD IID OOD IID OOD IID OOD IID OOD
UNICORN 1312.59±31.72 1301.17±25.98 -21.88±0.65 -93.57±4.85 266.85±14.54 235.55±17.69 315.90±5.67 304.45±10.88 407.83±75.58 392.48±74.45 2774.67±241.04 2603.68±182.85
CSRO 1180.48±228.10 458.08±253.00 -27.98±0.90 -102.44±5.17 276.18±19.37 233.33±12.37 309.63±6.50 301.12±9.99 310.47±58.29 278.94±65.37 2719.50±234.55 2800.93±182.06
CORRO 704.20±449.46 244.95±145.83 -37.11±2.60 -112.29±1.62 147.93±12.85 120.52±12.48 283.04±7.89 271.48±12.79 277.07±37.91 213.43±48.33 2467.83±174.57 2321.87±327.19
FOCAL 1186.03±271.58 860.58±252.96 -22.22±0.88 -96.72±1.69 217.10±29.14 173.16±24.37 301.84±4.48 296.87±12.76 308.05±98.42 286.34±91.08 2423.67±255.95 2315.56±303.03
Supervised 961.62±355.84 781.51±429.12 -23.81±0.71 -104.31±0.65 237.52±38.81 202.19±37.56 306.54±9.99 293.65±7.99 256.52±60.39 209.64±27.72 2489.37±248.05 2283.26±205.36
MACAW 1154.83±10.11 449.56±6.54 -56.04±2.42 -188.44±0.91 26.29±2.80 0.18±0.31 218.52±6.01 205.39±2.47 141.25±8.83 129.63±5.09 2430.67±157.42 1728.01±78.98
Prompt-DT 1176.54±39.64 -25.10±8.74 -118.14±65.59 -248.64±21.38 1.00±0.38 0.04±0.11 234.19±5.35 202.51±4.96 184.73±8.57 155.78±16.64 2164.86±85.20 1895.51±110.64
Figure 4 illustrates the learning curves of UNICORN vs. all baselines. IID entries of Table 1 depicts the performance of in-distribution generalization where the context for each task is collected by the behavior policies trained on the same task. The results demonstrate that UNICORN almost unanimously achieves SOTA performance across all benchmarks. The observation that FOCAL, CORRO, and CSRO achieve asymptotic performance close to UNICORN is also expected, since they are special cases of UNICORN according to our theory. Notably, UNICORN exhibits remarkable stability, especially on HalfCheetah-Dir and Walker-Param, where the other context-based methods suffer a significant performance decline during the later training process. Although the gradient-based MACAW and the transformer-based Prompt-DT show faster convergence, their asymptotic performance leaves much to be desired. Moreover, the average training time for MACAW is almost three times longer than the context-based methods.
4.4 Few-Shot Generalization to Out-of-Distribution Behavior Policies
Denote by $\{\pi_{\beta}^{i,t}\}_{i,t}$ the checkpoints of behavior policies for all tasks logged during data collection (i.e., training SAC agents), where $i$ refers to the task index and $t$ represents training iteration. For each test task $M^{j}$, we use every individual behavior policy in $\{\pi_{\beta}^{i,t}\}_{i,t}$ to collect the OOD contexts and the conditioned rollouts, as visualized in Figure 1. The OOD testing performance is measured by averaging returns across all testing tasks, as shown in the OOD entries of Table 1. While all methods suffer notable decline when facing OOD context, UNICORN maintains the strongest performance by a large margin. Another observation is that the context-based methods are in general significantly more robust compared to the gradient-based MACAW and transformer-based Prompt-DT, which suggests that context-based paradigm should be prioritized for research of generalizable OMRL. This also delineates the storyline of COMRL development from 2021 to date
$\underset{(2021)}{\text{FOCAL}}\rightarrow\underset{(2022)}{\text{CORRO}}\rightarrow\underset{(2023)}{\text{CSRO}}\rightarrow\underset{(2024)}{\text{UNICORN}}$
as a roadmap for pursuing more robust and generalizable task representation learning for COMRL.
4.5 Influence of Data Quality
Table 2: UNICORN vs. baselines on Ant-Dir datasets of various qualities. Each result is averaged by 6 random seeds. The best is bolded and the second best is underlined.
Algorithm Random Medium Expert
IID OOD IID OOD IID OOD
UNICORN 81.31±17.80 62.20±6.15 219.86±23.04 242.74±10.40 278.82±10.46 261.83±13.49
CSRO 1.55±2.65 -0.20±0.52 165.65±9.89 198.15±17.35 251.77±39.16 201.57±45.29
CORRO 0.87±0.53 -0.05±0.06 8.16±5.02 -7.01±2.37 -3.90±9.59 -13.59±9.06
FOCAL 67.00±26.33 43.65±10.08 171.10±84.26 187.06±86.45 229.02±42.07 246.27±20.21
Supervised 64.97±5.57 46.90±11.86 149.21±50.21 109.85±80.09 249.11±33.47 214.55±59.56
MACAW 3.412±0.74 -0.27±0.25 27.69±2.17 1.34±0.75 87.86±42.69 0.64±1.49
Prompt-DT 0.54±0.29 0.02±0.09 2.14±4.36 0.33±0.57 77.94±14.67 1.24±2.18
To test whether the UNICORN framework can be applied to different data distributions, we collect three types of datasets whose size is equal to the mixed dataset used in Section 4.3 and 4.4: random, medium and expert, which are characterized by the quality of the behavior policies. Each dataset is collected by one SAC agent. For random datasets, the SAC agent is randomly initialized and untrained.
As shown in Table 2, UNICORN demonstrates SOTA performance on all types of datasets. Notably, despite CSRO having better IID results than UNICORN on Ant-Dir under the mixed data distribution, UNICORN surpasses CSRO by a large margin under the narrow data distribution (medium and expert) and random data distribution.
We observe that CORRO fails on all three datasets. This may be due to the reliance of CORRO on the negative sample generator. Under narrow data distribution or poor random data distribution, it might be more challenging to generate negative samples of good quality. As for Prompt-DT, its performance improves significantly with increased data quality, which is expected since the decision transformer adopts a behavior-cloning-like supervised learning style.
5 Discussion
This section presents more empirical evidence on the broader applicability of the UNICORN framework.
5.1 Is UNICORN Model-Agnostic?
Since UNICORN tackles task representation learning from a general information theoretic perspective, it is in principle model-agnostic. Hence a straightforward idea is to transfer UNICORN to other model architectures like DT, which we envision to be a promising backbone for large-scale training of RL foundation models.
We employ a simple implementation by embedding the task representation vector $\bm{z}$ as the first token in sequence to prompt the learning and generation of DT. Instead of the unlearnable prompt used in the original Prompt-DT, we enforce task representation learning of the prompt by optimizing $\mathcal{L}_{\text{UNICORN}}$ and $\mathcal{L}_{\text{FOCAL}}$. We name these DT implementations as UNICORN-DT and FOCAL-DT respectively.
As shown in Table 3 and Figure 8, applying UNICORN and FOCAL on DT results in significant improvement in both IID and OOD generalization, with UNICORN being the superior option. Despite the gap between our implementation of UNICORN-DT and its MLP counterpart in terms of asymptotic performance, we expect UNICORN-DT to extrapolate favorably in the regime of greater dataset size and model parameters due to the scaling law of transformer (Kaplan et al., 2020) and DT (Lee et al., 2022). We leave this verification to future work.
Table 3: DT implementation of COMRL on HalfCheetah-Dir and Hopper-Param. Each result is averaged by 6 random seeds.
Algorithm HalfCheetah-Dir Hopper-Param
IID OOD IID OOD
UNICORN 1312.59±31.72 1301.17±25.98 315.90±5.67 304.45±10.88
UNICORN-DT 1233.19±10.42 1185.76±42.71 303.66±3.51 290.88±3.68
FOCAL-DT 1208.58±33.38 652.14±35.86 293.25±4.17 283.86±4.81
Prompt-DT 1176.54±39.64 -25.10±8.74 234.19±5.35 202.51±4.96
5.2 Can UNICORN be Exploited for Model-Based Paradigms?
As alluded in Section 3.3, UNICORN can potentially enable a model-based paradigm by generating data using the decoder $p_{\theta}(\bm{x}_{t}|\bm{z},\bm{x}_{b})$ and customized latent $\bm{z}$, which can be especially useful for generalization to OOD tasks. We implement this idea on Ant-Dir by sorting the tasks according to the polar angle of their goal direction ($\theta\in[0,2\pi)$), and construct task-level OOD by taking 20 consecutive tasks in the middle for training and the rest for testing (Figure 1). Gaussian noise $\bm{\epsilon}$ is applied to the real task representation $\bm{z}$ and $(\bm{s},\bm{a},\bm{z}+\bm{\epsilon})$ is then fed to the decoder to generate $(\bm{s}^{\prime},r)$. The imaginary rollouts $\{(\bm{s}_{t},\bm{a}_{t},\bm{s}^{\prime}_{t},r_{t})\}_{t=1}^{n}$ are used during the training of RL agent. We adopt the conventional ensemble technique in model-based RL (Yu et al., 2020b; Zhang et al., 2023) to stabilize the training process.
As shown in Figure 5, UNICORN-enabled model-based RL is the only method to achieve positive performance in this extremely challenging scenario where context shift is induced by OOD tasks.
Figure 5: Testing returns for OOD tasks. The learning curve is averaged by 6 random seeds.
6 Conclusion
In this paper, we unify three major context-based offline meta-RL algorithms, FOCAL, CORRO and CSRO (potentially a lot more) into a single framework based on the information bottleneck principle. Our theory offers valuable insight on how these methods are inherently connected and incrementally evolved in pursuit of better generalization against context shift. Based on the proposed framework, we arrive at a novel algorithm called UNICORN which is demonstrated to be remarkably robust and broadly applicable through extensive experiments. We believe our study offers potential for new implementations, optimality bounds and algorithms for both fully offline and offline-to-online COMRL paradigms. We leave these directions for future work.
Appendix A Pseudo-Code
Algorithm 1 UNICORN Meta-training
0:  Offline Datasets $\bm{X}=\{X^{i}\}_{i=1}^{N_{train}}$, training tasks $\bm{M}=\{M^{i}\}_{i=1}^{N_{train}}$, initialized learned policy $\pi_{\omega}$, Q function $Q_{\psi}$, context encoder $q_{\phi}$, decoder $p_{\theta}$, hyper-parameter $\alpha$, $\beta$, task batch size $N_{tbatch}$
1:  while not done do
2:     Sample task batch $\{M^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{M}$ and the corresponding replay buffer =$\mathcal{R}=\{X^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{X}$
3:     for step in each iter do
4:        // Train the context encoder and decoder
5:        Sample context $\{\bm{c}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
6:        Obtain task embeddings $\{\bm{z}^{j}\sim q_{\phi}(\bm{z}|\bm{c}^{j})\}_{j=1}^{N_{tbatch}}$
7:        Compute $\mathcal{L}_{\text{UNICORN}}=\mathcal{L}_{\text{recon}}+\frac{\alpha\beta-1}{(1-\alpha)\beta}\mathcal{L}_{\text{FOCAL}}$
8:        Update $\phi,\theta$ to minimize $\mathcal{L}_{\text{UNICORN}}$
9:        // Train the actor and critic
10:        Detach the task embeddings $\{\bm{z}^{j}\}_{j=1}^{N_{tbatch}}$
11:        Sample training data $\{\bm{d}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
12:        Compute $\mathcal{L}_{\text{actor}},\mathcal{L}_{\text{critic}}$
13:        Update $\omega,\psi$ to minimize $\mathcal{L}_{\text{actor}},\mathcal{L}_{\text{critic}}$
14:     end for
15:  end while
Algorithm 2 UNICORN Meta-testing
0:  Offline Datasets $\bm{X}=\{X^{i}\}_{i=1}^{N_{test}}$, testing tasks $\bm{M}=\{M^{i}\}_{i=1}^{N_{test}}$, learned policy $\pi_{\omega}$, context encoder $q_{\phi}$
1:  for each $M^{i}$ do
2:     Sample context $\bm{c}^{i}\sim X^{i}$
3:     Obtain task embedding $\bm{z}^{i}\sim q_{\phi}(\bm{z}|\bm{c}^{i})$
4:     Rollout policy $\pi_{\omega}(\bm{a}|\bm{s},\bm{z}^{i})$ for evaluation
5:  end for
Appendix B Proofs
Theorem B.1.
Let $\equiv$ denote equality up to a constant,
$\underbrace{I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})}_{\text{primary causality}}\quad\leq\quad I(\bm{Z};\bm{M})\quad\leq\underbrace{I(\bm{Z};\bm{X})}_{\text{primary + lesser causality}}$
holds up to a constant, where
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
[…]
Proof.
Assuming 2, $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})\equiv-\mathcal{L}_{\text{CORRO}}\leq I(\bm{Z};\bm{M})$ is shown in the Theorem 4.1. of CORRO [Yuan and Lu, 2022] and GENTLE [Zhou et al., 2023] with strong assumptions. We hereby present a proof with minimal assumptions. Given $I(\bm{Z};\bm{M};\bm{X}_{b})\geq 0$ in Assumption 3.2, we have
$I(\bm{Z};\bm{M})$ $=I(\bm{Z};\bm{M}|\bm{X}_{b})+I(\bm{Z};\bm{M};\bm{X}_{b})$
$\geq I(\bm{Z};\bm{M}|\bm{X}_{b})$
$=I(\bm{M};\bm{Z},\bm{X}_{t}|\bm{X}_{b})-I(\bm{M};\bm{X}_{t}|\bm{Z},\bm{X}_{b})$
$=\underbrace{I(\bm{Z};\bm{M}|\bm{X}_{t},\bm{X}_{b})}_{0\text{ by }\bm{M}\rightarrow\bm{X}\rightarrow\bm{Z}}+I(\bm{M};\bm{X}_{t}|\bm{X}_{b})-I(\bm{M};\bm{X}_{t}|\bm{Z},\bm{X}_{b})$
$=I(\bm{M};\bm{X}_{t}|\bm{X}_{b})-I(\bm{M};\bm{X}_{t}|\bm{Z},\bm{X}_{b})$
$=I(\bm{M};\bm{X}_{t}|\bm{X}_{b})-H(\bm{X}_{t}|\bm{Z},\bm{X}_{b})+H(\bm{X}_{t}|\bm{M},\bm{Z},\bm{X}_{b})$
$\geq I(\bm{M};\bm{X}_{t}|\bm{X}_{b})-H(\bm{X}_{t}|\bm{Z},\bm{X}_{b})$
$=\underbrace{I(\bm{M};\bm{X}_{t}|\bm{X}_{b})}_{\text{const}}-\underbrace{H(\bm{X}_{t})}_{\text{const}}+I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$
$\equiv I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$
$=I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+\underbrace{I(\bm{X}_{t};\bm{X}_{b})}_{\text{const}}$
$\equiv I(\bm{Z};\bm{X}_{t}|\bm{X}_{b}),$
as desired. The third and fourth lines are obtained by direct application of the chain rule of mutual information [Yeung, 2008]. All mutual information terms without $\bm{Z}$ are held constant in the fully offline scenario.
$I(\bm{Z};\bm{M})\leq I(\bm{Z};\bm{X})$ is a direct consequence of the Data Processing Inequality [Yeung, 2008] and the Markov chain $\bm{M}\rightarrow\bm{X}\rightarrow\bm{Z}$ in Assumption 3.2. We now prove the three claims regarding FOCAL, CORRO and CSRO:
  1. 1.
    $\mathcal{L}_{\text{FOCAL}}\equiv-I(\bm{Z};\bm{X})$.
    $I(\bm{Z};\bm{X})$ $:=\mathbb{E}_{\bm{x},\bm{z}}\left[\log\left(\frac{p(\bm{z},\bm{x})}{p(\bm{z})p(\bm{x})}\right)\right]$
    $=\mathbb{E}_{\bm{x},\bm{z}}\left[\log\left(\frac{1}{\frac{p(\bm{z})}{p(\bm{z}|\bm{x})}|\mathcal{M}|}\right)\right]+\log(|\mathcal{M}|)$
    $=\mathbb{E}_{\bm{x},\bm{z}}\left[\log\left(\frac{1}{\frac{p(\bm{z})}{p(\bm{z}|\bm{x})}|\mathcal{M}|\mathbb{E}_{\bm{x}}\left[\frac{p(\bm{z}|\bm{x})}{p(\bm{z})}\right]}\right)\right]+\log(|\mathcal{M}|)$
    $\approx\mathbb{E}_{\bm{x},\bm{z}}\left[\log\left(\frac{1}{\frac{p(\bm{z})}{p(\bm{z}|\bm{x})}\sum_{M^{i}\in\mathcal{M}}\mathbb{E}_{\bm{x}^{i}\sim X^{i}}\left[\frac{p(\bm{z}|\bm{x}^{i})}{p(\bm{z})}\right]}\right)\right]+\log(|\mathcal{M}|)$
    $=\mathbb{E}_{\bm{x},\bm{z}}\left[\log\left(\frac{\frac{p(\bm{z}|\bm{x})}{p(\bm{z})}}{\sum_{M^{*}\in\mathcal{M}}\mathbb{E}_{\bm{x}^{i}\sim X^{i}}\left[\frac{p(\bm{z}|\bm{x}^{i})}{p(\bm{z})}\right]}\right)\right]+\log(|\mathcal{M}|)$
    $=\mathbb{E}_{\bm{x},\bm{z}}\left[\log\left(\frac{h(\bm{x},\bm{z})}{\sum_{M^{i}\in\mathcal{M}}h(\bm{x}^{i},\bm{z})}\right)\right]+\log(|\mathcal{M}|)$
    The first term on RHS is precisely the supervised contrastive learning objective introduced by a variant of FOCAL [Li et al., 2021a], which is equivalent to the negative distance metric learning loss $\mathcal{L}_{\text{FOCAL}}$ with the effect of pushing away embeddings of different tasks while pulling together those from the same task. Therefore we have
    $I(\bm{Z};\bm{X})$ $=\mathbb{E}_{\bm{x},\bm{z}}\left[\log\left(\frac{h(\bm{x},\bm{z})}{\sum_{M^{i}\in\mathcal{M}}h(\bm{x}^{i},\bm{z})}\right)\right]+\text{const}$
    $\equiv-\mathcal{L}_{\text{FOCAL}}.$
  2. 2.
    $\mathcal{L}_{\text{CORRO}}\equiv-I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$.
    $I(\bm{Z};\bm{X_{t}}|\bm{X_{b}})$ $:=\mathbb{E}_{\bm{x}_{t},\bm{x}_{b},\bm{z}}\left[\log\left(\frac{p(\bm{z},\bm{x_{t}}|\bm{x_{b}})}{p(\bm{z}|\bm{\bm{x_{b}}})p(\bm{x_{t}}|\bm{\bm{x_{b}}})}\right)\right]$
    $=\mathbb{E}_{\bm{x_{t}},\bm{\bm{x_{b}}},\bm{z}}\left[\log\left(\frac{1}{\frac{p(\bm{z}|\bm{\bm{x_{b}}})}{p(\bm{z}|\bm{x_{t}},\bm{\bm{x_{b}}})}|\mathcal{M}|}\right)\right]+\log(|\mathcal{M}|)$
    $=\mathbb{E}_{\bm{x_{t}},\bm{\bm{x_{b}}},\bm{z}}\left[\log\left(\frac{1}{\frac{p(\bm{z}|\bm{x}_{b})}{p(\bm{z}|\bm{X})}|\mathcal{M}|\mathbb{E}_{M^{*}\in\mathcal{M}}\left[\frac{p(\bm{z}|\bm{x}^{*}_{t},\bm{\bm{x_{b}}})}{p(\bm{z}|\bm{\bm{x_{b}}})}\right]}\right)\right]+\log(|\mathcal{M}|)$
    $\approx\mathbb{E}_{\bm{x_{t}},\bm{\bm{x_{b}}},\bm{z}}\left[\log\left(\frac{1}{\frac{p(\bm{z}|\bm{\bm{x_{b}}})}{p(\bm{z}|\bm{x_{t}},\bm{\bm{x_{b}}})}\sum_{M^{*}\in\mathcal{M}}\frac{p(\bm{z}|\bm{x}^{*}_{t},\bm{\bm{x_{b}}})}{p(\bm{z}|\bm{\bm{x_{b}}})}}\right)\right]+\log(|\mathcal{M}|)$
    $=\mathbb{E}_{\bm{x_{t}},\bm{\bm{x_{b}}},\bm{z}}\left[\log\left(\frac{\frac{p(\bm{z}|\bm{x_{t}},\bm{\bm{x_{b}}})}{p(\bm{z}|\bm{\bm{x_{b}}})}}{\sum_{M^{*}\in\mathcal{M}}\frac{p(\bm{z}|\bm{x}^{*}_{t},\bm{\bm{x_{b}}})}{p(\bm{z}|\bm{\bm{x_{b}}})}}\right)\right]+\log(|\mathcal{M}|)$
    $=\mathbb{E}_{\bm{x_{t}},\bm{\bm{x_{b}}},\bm{z}}\left[\log\left(\frac{\frac{p(\bm{z}|\bm{x_{t}},\bm{\bm{x_{b}}})}{p(z)}}{\sum_{M^{*}\in\mathcal{M}}\frac{p(\bm{z}|\bm{x}^{*}_{t},\bm{\bm{x_{b}}})}{p(z)}}\right)\right]+\log(|\mathcal{M}|)$
    $=\mathbb{E}_{\bm{x},\bm{z}}\left[\log\left(\frac{h(\bm{x},\bm{z})}{\sum_{M^{*}\in\mathcal{M}}h(\bm{x}^{*},\bm{z})}\right)\right]+\log(|\mathcal{M}|)$
    $\equiv-\mathcal{L}_{\text{CORRO}}.$
  3. 3.
    $\mathcal{L}_{\text{CSRO}}\geq(\lambda-1)I(\bm{Z};\bm{X})-\lambda I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$.
    The CLUB loss in Eqn 6 operates as a upper bound of $I(\bm{Z};\bm{X}_{b})$ [Cheng et al., 2020]. By Eqn 5 we have
    $\mathcal{L}_{\text{CSRO}}$ $=\mathcal{L}_{\text{FOCAL}}+\lambda L_{\text{CLUB}}$
    $\geq\mathcal{L}_{\text{FOCAL}}+\lambda I(\bm{Z};\bm{X}_{b})$
    $=\mathcal{L}_{\text{FOCAL}}+\lambda\underbrace{(I(\bm{Z};\bm{X})-I(\bm{Z};\bm{X}_{t}|\bm{X}_{b}))}_{\text{chain rule of mutual information}}$
[…]
    $=(\lambda-1)I(\bm{Z};\bm{X})-\lambda I(\bm{Z};\bm{X}_{t}|\bm{X}_{b}).$
Appendix C Further Experiments
Figure 6: Hyper-parameter study results in all environments, each result is averaged over 6 random seeds.
C.1 Hyper-parameter Study of $\frac{\alpha\beta-1}{(1-\alpha)\beta}$
We experiment with $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ from 0 to 10, as shown in Figure 6. We observe that UNICORN has an optimal $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ in all environments. This observation suggests that it is insufficient to use either FOCAL loss or reconstruction loss alone, emphasizing the necessity of their combination proposed in Eqn 7 and 8.
Intuitively, for tasks differ in transition dynamics, the lesser causality $I(\bm{Z};\bm{X}_{b})$ should be higher since the state distribution $\mu_{\pi}(\bm{s})$ vary across tasks even with the same behavior policy $\pi$. This indicates that the optimal proportion of FOCAL loss in Eqn 12 should be greater on these environments, as it maximizes $I(\bm{Z};\bm{X})$ along with the primary causality $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$. Such thought experiment is verified in this study. We record the optimal $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ of all benchmarks in Table 5, demonstrating that for tasks differ in reward function (Ant-Dir, HalfCheetah-Dir) and transition dynamics (Hopper-Param, Walker-Param), the optimal value of $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ is around 0.1 and 1 respectively. The fact that the optimal $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ does not vanish for Ant-Dir and HalfCheetah-Dir suggests that the lesser causality $I(\bm{Z};\bm{X}_{b})$ should be nonzero even when all tasks share the same transition dynamics. We conjecture this is partially due to the fact that the behavior policy we use for each task is also trained on that task, which leads to task-dependent state distribution $\mu_{\pi}(\bm{s})$. However, this applies to the IID setting only.
C.2 UNICORN-0: A Label-free Version
When $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ equals zero, UNICORN task representation learning reduces to the minimization of the reconstruction loss only, which becomes a label-free algorithm (i.e., it does not require the knowledge of task identities/labels to optimize, as opposed to the contrastive learning in previous methods). We name this special case UNICORN-0, which is equivalent to a concurrent method GENTLE [Zhou et al., 2023]. For a fair comparison, we use a variant of a Bayesian OMRL method BOReL [Dorfman et al., 2021] that does not utilize oracle reward functions, as a label-free baseline. At test time, we modify BOReL to use offline datasets rather than the online data collected by interacting with the environment to infer the task information [^1] Experiment shows that the usage of offline dataset rather than online data has little effect on the results of BOReL.. As shown in Table 4, UNICORN-0 is also competitive with BOReL.
Table 4: UNICORN-0 compared to another label-free COMRL method, BOReL, on Ant-Dir.
Algorithm Ant-Dir
IID OOD
UNICORN-0 219.88±16.24 199.86±9.26
BOReL 206.4±17.93 186.55±9.54
C.3 Visualization of Task Representations
For better interpretation, we visualize the task representations by 2D projection of the embedding vectors via t-SNE [Van der Maaten and Hinton, 2008]. For each test task, we generate a trajectory using each policy in $\{\pi_{\beta}^{i,t}\}_{i,t}$ and infer their task representations $\bm{z}$. This setting aligns with our OOD experiments in the main text.
Since the UNICORN objective operates as a convex combination of the reconstruction loss (UNICORN-0) and the FOCAL loss, we compare representations of UNICORN to these two extremes. As shown in Figure 7, compared to UNICORN-0, UNICORN provides evidently more distinguishable task representations for OOD data, emphasizing the necessity of joint optimization with the FOCAL loss. On the other hand, despite the seemingly higher quality of task representations of FOCAL, the performance of FOCAL is much worse than UNICORN. We speculate that since FOCAL blindly separates embeddings from different tasks, it may fail to capture shared structure between similar tasks, whereas UNICORN is able to do so by generative modeling via the reconstruction loss.
Figure 7: The 2D projection of the learned task representation space in Ant-Dir. Points are uniformly sampled from out-of-distribution data. Tasks of given goals from 0 to 6 are mapped to rainbow colors, ranging from purple to red.
C.4 More on UNICORN-DT
We provide the learning curves of the results in Section 5.1 here. As shown in Figure 8, UNICORN-DT demonstrates both the higher convergence speed and the higher asymptotic performance compared to the vanilla Prompt-DT.
Figure 8: Test returns of UNICORN-DT against UNICORN and Prompt-DT on 2 benchmarks. The learning curve is averaged by 6 random seeds.
Appendix D Experimental Details
Table 5 lists the necessary hyper-parameters that we used to produce the experimental results.
Table 5: Configurations and hyper-parameters used in the training process.
Configurations Ant-Dir HalfCheetah-Dir HalfCheetah-Vel Hopper-Param Walker-Param Reach
dataset size 1e5 2e5 2e5 3e5 4.5e5 1e5
task representation dimension 5 5 5 40 32 2
optimal $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ 0.15 0.01 0.15 1.5 0.7 0.3
training steps 200k
task batch size 16
RL batch size 256
context training size 200 500
learning rate 3e-4
RL network width 256
RL network depth 3
encoder width 64 128 64

### v2

@@
 Among which, context-based OMRL (COMRL) as a popular paradigm, aims to learn a universal policy conditioned on effective task representations.
-In this work, by examining several key milestones in the field of COMRL, we propose to integrate these seemingly independent methodologies into a unified information theoretic framework.
+In this work, by examining several key milestones in the field of COMRL, we propose to integrate these seemingly independent methodologies into a unified framework.
 Most importantly, we show that the pre-existing COMRL algorithms are essentially optimizing the same mutual information objective between the task variable $\bm{M}$ and its latent representation $\bm{Z}$ by implementing various approximate bounds.
-Based on the theoretical insight and the information bottleneck principle, we arrive at a novel algorithm dubbed UNICORN, which exhibits remarkable generalization across a broad spectrum of RL benchmarks, context shift scenarios, data qualities and deep learning architectures, attaining the new state-of-the-art.
-We believe that our framework could open up avenues for new optimality bounds and COMRL algorithms.
+Such theoretical insight offers ample design freedom for novel algorithms.
+As demonstrations, we propose a supervised and a self-supervised implementation of $I(\bm{Z};\bm{M})$, and empirically show that the corresponding optimization algorithms exhibit remarkable generalization across a broad spectrum of RL benchmarks, context shift scenarios, data qualities and deep learning architectures.
+This work lays the information theoretic foundation for COMRL methods, leading to a better understanding of task representation learning in the context of reinforcement learning  [^1] The open-source code is at https://github.com/betray12138/UNICORN.git..
 1 Introduction
@@
 In order to quickly adapt to the unknown, the agents need to integrate prior knowledge with minimal experience (namely the context) collected from the new tasks or environments, without over-fitting to the new data.
-This meta-RL mechanism has been adopted in many applications such as games (Zhao et al., 2022a; Beck et al., 2023), robotics (Yu et al., 2020a; Kumar et al., 2021) and drug discovery (Feng et al., 2022).
-However, for data collection, classical meta-RL usually requires enormous online explorations of the environments (Finn et al., 2017; Rakelly et al., 2019), which is impractical in many safety-critical scenarios such as healthcare (Gottesman et al., 2019) and autonomous driving (Shalev-Shwartz et al., 2016).
-As a remedy, offline RL (Levine et al., 2020) enables agents to learn from logged experience only, therefore circumventing risky or costly online interactions.
-Recently, offline meta-RL (OMRL) (Li et al., 2021b; Mitchell et al., 2021; Dorfman et al., 2021) has emerged as a novel paradigm to significantly extend the applicable range of RL by "killing two birds in one stone": it builds powerful agents that can quickly learn and adapt by meta-learning, while leveraging offline RL mechanism to ensure a safe and efficient optimization procedure.
-In the context of supervised or self-supervised learning, which is de facto offline, OMRL is reminiscent of the multi-task learning (Raffel et al., 2020), meta-training (Finn et al., 2017) and fine-tuning (Hu et al., 2021; Li and Liang, 2021; Lester et al., 2021) of pre-trained large models.
-We envision it to be a cornerstone of RL foundation models (Bommasani et al., 2021; Reed et al., 2022) in the future.
+This meta-RL mechanism has been adopted in many applications such as games [1, 2], robotics [3, 4] and drug discovery [5].
+However, for data collection, classical meta-RL usually requires enormous online explorations of the environments [6, 7], which is impractical in many safety-critical scenarios such as healthcare [8] and robotic manipulation [9].
+As a remedy, offline RL [10] enables agents to learn from logged experience only, therefore circumventing risky or costly online interactions.
+Recently, offline meta-RL (OMRL) [11, 12, 13] has emerged as a novel paradigm to significantly extend the applicable range of RL by "killing two birds in one stone": it builds powerful agents that can quickly learn and adapt by meta-learning, while leveraging offline RL mechanism to ensure a safe and efficient optimization procedure.
+In the context of classical supervised or self-supervised learning, which is de facto offline, OMRL is reminiscent of the multi-task learning [14], meta-training [6] and fine-tuning [15, 16, 17] of pre-trained large models.
+We envision it as a cornerstone of RL foundation models  [18, 19, 20] in the future.
 Along the line of OMRL research, context-based offline meta-reinforcement learning (COMRL) is a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs).
-Intuitively, the crux of COMRL is learning effective task representations, and therefore enabling the agent to react optimally and adaptively in various contexts.
-To achieve this goal, one of the earliest COMRL algorithms FOCAL (Li et al., 2021b) proposes to capture the structure of task representations by distance metric learning.
-From a geometric perspective, it essentially performs clustering by pushing away latent embeddings of different tasks while pulling together those from the same task, hence ensuring consistent and distinguishable task representations.
-Despite its effectiveness, FOCAL is reported to be vulnerable to context shifts (Li et al., 2021a), i.e., when testing on out-of-distribution (OOD) data.
-Such problems are particularly challenging for OMRL, since any context shift incurred at test time can not be rectified in the fully offline setting, which may result in severely degraded generalization performance (Li et al., 2021a; Gao et al., 2023).
-To alleviate the problem, follow-up works such as CORRO (Yuan and Lu, 2022) reformulates the task representation learning of COMRL as maximizing the mutual information $I(\bm{Z};\bm{M})$ between the task variable $\bm{M}$ and its latent $\bm{Z}$.
-It then approximates $I(\bm{Z};\bm{M})$ by an InfoNCE (Oord et al., 2018) contrastive loss, where the positive and negative pairs are conditioned on identical state-action tuples $(\bm{s},\bm{a})$.
-Inspired by CORRO, a recently proposed method CSRO (Gao et al., 2023) introduces an additional mutual information term between $\bm{Z}$ and $(\bm{s},\bm{a})$.
-By explicitly minimizing it along with the FOCAL objective, CSRO is demonstrated to achieve the state-of-the-art (SOTA) generalization performance on various MuJoCo (Todorov et al., 2012) benchmarks.
-In this paper, following the recent development and storyline of COMRL, we present a novel and unified view of several signature methods including FOCAL, CORRO and CSRO under a information theoretic framework.
-We first show that these algorithms are essentially optimizing different approximate bounds of the same mutual information objective $I(\bm{Z},\bm{M})$.
-Then based on our proposed Unified Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning (UNICORN), we arrive at a new COMRL optimization scheme by deriving a unified and generalized task representation learning objective in light of the information bottleneck principle  (Tishby et al., 1999; Tishby and Zaslavsky, 2015).
-We demonstrate experimentally that the proposed method achieves state-of-the-art in-distribution and OOD generalization performance on a wide range of RL domains, OOD settings, data qualities and model architectures.
-Our framework potentially provides a principled roadmap to new COMRL algorithms by seeking better optimality bounds/approximations of the objective, as well as new implementations to further combat context shift.
-2 Related Works
-Offline Meta-Reinforcement Learning.
-OMRL solves the problems of generalization and expensive data collection by meta-learning on pre-collected offline datasets.
-To overcome the overestimation of out-of-distribution actions (Kumar et al., 2019; Levine et al., 2020; Gao et al., 2022), offline RL usually leverages regularizations such as KL-divergence (Jaques et al., 2019), maximum mean discrepancy (MMD) (Kumar et al., 2019) and Wasserstein Distance (Gulrajani et al., 2017) to constrain the distribution mismatch between the behavior and learner policies.
-Moreover, a general mechanism called behavior regularized actor critic (BRAC) (Wu et al., 2019) involving value penalty and behavior regularization has proven successful in several previous studies (Li et al., 2021b; Gao et al., 2023), and is adopted in our method.
-From a meta-learning perspective, the realization of OMRL can be categorized as gradient-based and context-based paradigms.
-Gradient-based meta-RL methods (Finn et al., 2017; Mitchell et al., 2021) embed an optimization procedure such as gradient descent into the meta-level iteration loops, which are model-agnostic but tend to reach lower asymptotic performance with less computational efficiency in practice.
-COMRL conditions the models on a contextual representation of MDP (Hallak et al., 2015) generated by recurrent (Wang et al., 2016; Fakoor et al., 2020), recursive (Mishra et al., 2017; Xu et al., 2022), probabilistic (Rakelly et al., 2019; Zintgraf et al., 2019) or deterministic (Li et al., 2021b, a; Yuan and Lu, 2022) task inference modules.
-At test time, it leverages neural networks as universal value/policy function approximators (Schaul et al., 2015) to perform meta-adaptation by transfer learning of the task representations.
-Depending on the source of the testing data, fully offline (Li et al., 2021b, a; Yuan and Lu, 2022) and offline-to-online (Zhao et al., 2022b; Pong et al., 2022) COMRL have been investigated.
-Notwithstanding, both paradigms may suffer significant OOD as the data-collecting policies might drastically differ from the training sets.
-In this work, we restrict our attention to fully offline COMRL, and demonstrate the remarkable robustness of our proposed framework to context shift in this setting.
+Intuitively, the crux of COMRL lies in learning effective task representations, hence enabling the agent to react optimally and adaptively in various contexts.
+To this end, one of the earliest COMRL algorithms FOCAL [11] proposes to capture the structure of task representations by distance metric learning.
+From a geometric perspective, it essentially performs clustering by repelling latent embeddings of different tasks while pulling together those from the same task, therefore ensuring consistent and distinguishable task representations.
+Despite its effectiveness, FOCAL is reported to be vulnerable to context shifts [21], i.e., when testing on out-of-distribution (OOD) data (Fig.
+1).
+Such problems are particularly challenging for OMRL, since any context shift incurred at test time can not be rectified in the fully offline setting, which may result in severely degraded generalization performance  [21, 22].
+To alleviate the problem, follow-up works such as CORRO [23] reformulates the task representation learning of COMRL as maximizing the mutual information $I(\bm{Z};\bm{M})$ between the task variable $\bm{M}$ and its latent $\bm{Z}$.
+It then approximates $I(\bm{Z};\bm{M})$ by an InfoNCE [24] contrastive loss, where the positive and negative pairs are conditioned on the state-action tuples $(\bm{s},\bm{a})$.
+Inspired by CORRO, a recently proposed method CSRO [22] introduces an additional mutual information term between $\bm{Z}$ and $(\bm{s},\bm{a})$.
+By explicitly minimizing it along with the FOCAL objective, CSRO is demonstrated to achieve the state-of-the-art (SOTA) generalization performance on various MuJoCo [25] benchmarks.
+Contributions  In this paper, following the recent development and storyline of COMRL, we present a Unified Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning (UNICORN) encompassing pre-existing methods.
+We first prove that the objectives of FOCAL, CORRO and CSRO operate as the upper bound, lower bound of $I(\bm{Z};\bm{M})$ and their linear interpolation respectively, which provides a nontrivial theoretical unification of these methods.
+Second, by the aforementioned insight and an analysis of the COMRL causal structures, we shed light on how CORRO and CSRO improve context-shift robustness compared to their predecessors by trading off causal and spurious correlations between $\bm{Z}$ and input data $\bm{X}$.
+Lastly, by examining eight related meta-RL methods (Table 1) concerning their objectives and implementations, we highlight the potential design choices of novel algorithms offered by our framework.
+As examples, we investigate two instantiated algorithms, one supervised and the other self-supervised, and demonstrate experimentally that they achieve competitive in-distribution and exceptional OOD generalization performance on a wide range of RL domains, OOD settings, data qualities and model architectures.
+Our framework provides a principled roadmap to novel COMRL algorithms by seeking better approximations/regularizations of $I(\bm{Z};\bm{M})$, as well as new implementations to further combat context shift.
 2 Method
 Figure 1: Context shift of COMRL in Ant-Dir.
-Left: Given a task specified by a goal direction (dashed line), the RL agent is trained on data generated by a variety of behavior policies trained on the same task (red).
-At test time, however, the context might be collected by behavior policies trained on different tasks (blue), causing a context shift of OOD behavior policies (Section 4.4).
-Middle: Under OOD testing context, UNICORN (red) is more robust than baselines such as FOCAL (green) in terms of navigating the Ant robot along the right direction.
+Left: Given a task $M^{i}$ specified by a goal direction (dashed line), the RL agent is trained on data generated by a variety of behavior policies trained on the same task $M^{i}$ (red).
+At test time, however, the context might be collected by behavior policies trained on different tasks $\{M^{j}\}$ (blue), causing a context shift of OOD behavior policies (Section 3.3).
+Middle: Against OOD context, UNICORN (red) is more robust than baselines such as FOCAL (green) in terms of navigating the Ant robot towards the right direction.
 Right: Besides behavior policy, the task distribution (e.g., goal positions in Ant) can induce significant context shift (Section 4.2), which is also a challenging scenario for COMRL models to generalize.
-3.1 Preliminaries and Problem Statement
-We consider MDP modeled as $M=(\mathcal{S},\mathcal{A},T,R,\mathcal{\rho}_{0},\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $T(\bm{s}^{\prime}|\bm{s},\bm{a})$, bounded reward function $R(\bm{s},\bm{a})$, initial state distribution $\rho_{0}(\bm{s})$ and discount factor $\gamma\in(0,1)$.
+2.1 Preliminaries, Problem Statement and Related Work
+We consider MDP modeled as $M=(\mathcal{S},\mathcal{A},T,R,\mathcal{\rho}_{0},\gamma,H)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $T(\bm{s}^{\prime}|\bm{s},\bm{a})$, bounded reward function $R(\bm{s},\bm{a})$, initial state distribution $\rho_{0}(\bm{s})$, discount factor $\gamma\in(0,1)$ and $H$ the horizon.
 The goal is to find a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ to maximize the expected return.
@@
 The V-function and Q-function are given by
-$V_{\pi}(\bm{s})$ $=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\bm{s}_{t}\sim\mu_{\pi}^{t}(\bm{s}),\bm{a}_{t}\sim\pi}[R(\bm{s}_{t},\bm{a}_{t})]$ (1)
+$V_{\pi}(\bm{s})$ $=\sum_{t=0}^{H-1}\gamma^{t}\mathbb{E}_{\bm{s}_{t}\sim\mu_{\pi}^{t}(\bm{s}),\bm{a}_{t}\sim\pi}[R(\bm{s}_{t},\bm{a}_{t})],$ (1)
 $Q_{\pi}(\bm{s},\bm{a})$ $=R(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s}^{\prime}\sim T(\bm{s}^{\prime}|\bm{s},\bm{a})}[V_{\pi}(\bm{s}^{\prime})].$ (2)
-In this paper, we restrict to scenarios where tasks share the same state and action space while differing in reward and transition functions, with each OMRL task defined as an instance of MDP: $M^{i}=(\mathcal{S},\mathcal{A},T^{i},R^{i},\rho_{0},\gamma)\in\mathcal{M}$, where $\mathcal{M}$ is the set of all possible MDPs to be considered.
+In this paper, we restrict to scenarios where tasks share the same state and action space and can only be distinguished via reward and transition functions.
+Each OMRL task is defined as an instance of MDP: $M^{i}=(\mathcal{S},\mathcal{A},T^{i},R^{i},\rho_{0},\gamma,H)\in\mathcal{M}$, where $\mathcal{M}$ is the set of all possible MDPs to be considered.
 For each task index $i$, the offline dataset $X^{i}=\{(\bm{s}_{j}^{i},\bm{a}_{j}^{i},r_{j}^{i},\bm{s}{{}^{\prime}}_{j}^{i})\}$ is collected by a behavior policy $\pi^{i}_{\beta}$.
 For meta-learning, given a trajectory segment $\bm{c}^{i}_{1:n}=\{(\bm{s}_{j}^{i},\bm{a}_{j}^{i},r_{j}^{i},\bm{s}{{}^{\prime}}_{j}^{i})\}_{j=1}^{n}$ of length $n$ as the context of $M^{i}$, COMRL employs a context encoder $q_{\bm{\phi}}(\bm{z}|\bm{c}_{1:n}^{i})$ to obtain the latent representation $\bm{z}^{i}$ of task $M^{i}$.
-If $\bm{z}$ contains enough information about the task identity, COMRL can be treated as a special case of RL on partially-observed MDP (Kaelbling et al., 1998), where $\bm{z}$ is interpreted as a faithful representation of the unobserved state.
-By conditioning on $\bm{z}$, the learning of universal policy $\pi_{\bm{\theta}}(\cdot|\bm{s},\bm{z})$ and value function $V_{\pi}(\bm{s},\bm{z})$ (Schaul et al., 2015) become regular RL, and optimality can be attained by Bellman updates with theoretical guarantees.
-To this end, FOCAL (Li et al., 2021b) proposes to decouple COMRL into the upstream task representation learning and downstream offline RL.
-For the former, FOCAL effectively clusters task embeddings by metric learning:
-$\mathcal{L}_{\text{FOCAL}}=$ $\,\,\underset{\bm{\phi}}{\min}\,\,\mathbb{E}_{i,j}\text{\LARGE\{}\textbf{1}\{i=j\}||\bm{z}^{i}-\bm{z}^{j}||_{2}^{2}+\textbf{1}\{i\neq j\}\frac{\beta}{||\bm{z}^{i}-\bm{z}^{j}||_{2}^{n}+\epsilon}\text{\LARGE\}}.$ (3)
-However, it is empirically shown that FOCAL struggles to generalize in the presence of context shift (Li et al., 2021a).
+If $\bm{z}$ contains sufficient information about the task identity, COMRL can be treated as a special case of RL on partially-observed MDP [26], where $\bm{z}$ is interpreted as a faithful representation of the unobserved state.
+By conditioning on $\bm{z}$, the learning of universal policy $\pi_{\bm{\theta}}(\cdot|\bm{s},\bm{z})$ and value function $V_{\pi}(\bm{s},\bm{z})$ [27] become regular RL, and optimality can be attained by Bellman updates with theoretical guarantees.
+To this end, FOCAL [11] proposes to decouple COMRL into the upstream task representation learning and downstream offline RL.
+For the former, which can be treated independently, FOCAL employs metric learning to achieve effective clustering of task embeddings:
+$\mathcal{L}_{\text{FOCAL}}=$ $\,\,\underset{\bm{\phi}}{\min}\,\,\mathbb{E}_{i,j}\left\{\mathbbm{1}\{i=j\}||\bm{z}^{i}-\bm{z}^{j}||_{2}^{2}+\mathbbm{1}\{i\neq j\}\frac{\beta}{||\bm{z}^{i}-\bm{z}^{j}||_{2}^{n}+\epsilon}\right\}.$ (3)
+However, it is empirically shown that FOCAL struggles to generalize in the presence of context shift  [21].
 This problem has been formulated in several earliest studies of OMRL.
 Li et al.
-(2020) observed that when there are large divergence between the state-action distributions of the offline datasets, due to shortcut learning (Geirhos et al., 2020), the task encoder may learn to ignore the causal information like rewards and spuriously correlate primarily state-action pairs to the task identity, leading to poor generalization performance.
+[28] observed that when there are large divergence of the state-action distributions among the offline datasets, due to shortcut learning [29], the task encoder may learn to ignore the causal information like rewards and spuriously correlate primarily state-action pairs to the task identity, leading to poor generalization performance.
 Dorfman et al.
-(2021) concurrently identified a related problem which they termed MDP ambiguity in the context of Bayesian offline RL.
+[13] concurrently identified a related problem which they termed MDP ambiguity in the context of Bayesian offline RL.
 An example is illustrated in Figure 1.
 The problem is further exacerbated in the fully offline setting, as the testing distribution is fixed and cannot be augmented by online exploration, allowing no theoretical guarantee for the context shift.
-To mitigate context shift, a subsequent algorithm CORRO (Yuan and Lu, 2022) proposes to optimize a lower bound of $I(\bm{Z};\bm{M})$ in the form of an InfoNCE (Oord et al., 2018) contrastive loss
+To mitigate context shift, a subsequent algorithm CORRO [23] proposes to optimize a lower bound of $I(\bm{Z};\bm{M})$ in the form of an InfoNCE [24] contrastive loss
 $\mathcal{L}_{\text{CORRO}}$ $=\underset{\bm{\phi}}{\min}\,\,\mathbb{E}_{\bm{x},\bm{z}}\left[-\log\left(\frac{h(\bm{x},\bm{z})}{\sum_{M^{*}\in\mathcal{M}}h(\bm{x}^{*},\bm{z})}\right)\right],$ (4)
-where $\bm{x}=(\bm{s},\bm{a},\bm{r},\bm{s}^{\prime})$, $\bm{z}\sim q_{\bm{\phi}}(\bm{z}|\bm{x})$, $h(\bm{x},\bm{z})=\frac{P(\bm{z}|\bm{x})}{P(\bm{z})}\approx\frac{q_{\bm{\phi}}(\bm{z}|\bm{x})}{P(\bm{z})}$, and $\bm{x}^{*}=(\bm{s},\bm{a},r^{*},\bm{s}^{\prime*})$ as a transition tuple generated for task $M^{*}\in\mathcal{M}$ conditioned on the same $(\bm{s},\bm{a})$.
-To further combat the spurious correlation between the task representation and behavior-policy-induced state-actions, a recently proposed method CSRO (Gao et al., 2023) introduces a CLUB upper bound (Cheng et al., 2020) of the mutual information between $\bm{z}$ and $(\bm{s},\bm{a})$, to regularize the FOCAL objective:
-$\mathcal{L}_{\text{CSRO}}$ $=\underset{\bm{\phi}}{\min}\left\{\mathcal{L}_{\text{FOCAL}}+\lambda L_{\text{CLUB}}\right\}$ (5)
+where $\bm{x}=(\bm{s},\bm{a},r,\bm{s}^{\prime})$, $\bm{z}\sim q_{\bm{\phi}}(\bm{z}|\bm{x})$, $h(\bm{x},\bm{z})=\frac{P(\bm{z}|\bm{x})}{P(\bm{z})}\approx\frac{q_{\bm{\phi}}(\bm{z}|\bm{x})}{P(\bm{z})}$, and $\bm{x}^{*}=(\bm{s},\bm{a},r^{*},\bm{s}^{\prime*})$ as a transition tuple generated for task $M^{*}\in\mathcal{M}$ conditioned on the same $(\bm{s},\bm{a})$.
+To further combat the spurious correlation between the task representation and behavior-policy-induced state-actions, a recently proposed method CSRO [22] introduces a CLUB upper bound [30] of the mutual information between $\bm{z}$ and $(\bm{s},\bm{a})$, to regularize the FOCAL objective:
+$\mathcal{L}_{\text{CSRO}}$ $=\underset{\bm{\phi}}{\min}\left\{\mathcal{L}_{\text{FOCAL}}+\lambda L_{\text{CLUB}}\right\},$ (5)
 $L_{\text{CLUB}}$ $=\mathbb{E}_{i}\left[\log q_{\bm{\phi}}(\bm{z}_{i}|\bm{s}_{i},\bm{a}_{i})-\mathbb{E}_{j}\left[\log q_{\bm{\phi}}(\bm{z}_{j}|\bm{s}_{i},\bm{a}_{i})\right]\right].$ (6)
-In the next section, we will show how these algorithms are inherently connected and can be further improved under our proposed unified information theoretic framework.
+In the next section, we will show how these algorithms are inherently connected and how their context-shift robustness gets improved incrementally.
 2.2 A Unified Information Theoretic Framework
@@
 Definition 2.1.
-Given an input context variable $\bm{X}\in\mathcal{X}$ and its associated task/MDP random variable $\bm{M}\in\mathcal{M}$, task representation learning in COMRL aims to find a minimal sufficient statistics $\bm{Z}$ of $\bm{X}$ with respect $\bm{M}$.
-In pure statistical terms, Definition 3.1 implies that an ideal representation $\bm{Z}$ is the simplest mapping of $\bm{X}$ that captures the mutual information $I(\bm{X};\bm{M})$.
-We therefore assume the following dependency structures in terms of directed graphical models:
-Assumption 3.2.
-The dependency graphs of COMRL are given by
-satisfying $I(\bm{Z};\bm{M};\bm{X}_{b})\geq 0$, where $\bm{X}_{b}$ and $\bm{X}_{t}$ are the behavior-related $(\bm{s},\bm{a})$-component and task-related $(\bm{s}^{\prime},r)$-component of the context $\bm{X}$, with $\bm{X}=(\bm{X}_{t},\bm{X}_{b})$.
+Given an input context variable $\bm{X}\in\mathcal{X}$ and its associated task/MDP random variable $\bm{M}\in\mathcal{M}$, task representation learning in COMRL aims to find a sufficient statistics $\bm{Z}$ of $\bm{X}$ with respect $\bm{M}$.
+In pure statistical terms, Definition 2.1 implies that an ideal representation $\bm{Z}$ is a mapping of $\bm{X}$ that captures the mutual information $I(\bm{X};\bm{M})$.
+We therefore define the following dependency structures in terms of directed graphical models:
+Definition 2.2.
+The dependency graphs of COMRL is given by Fig.
+2, where $\bm{X}_{b}$ and $\bm{X}_{t}$ are the behavior-related $(\bm{s},\bm{a})$-component and task-related $(\bm{s}^{\prime},r)$-component of the context $\bm{X}$, with $\bm{X}=(\bm{X}_{t},\bm{X}_{b})$.
+Figure 2: Graphical Models of COMRL.
 For the first graph, $\bm{M}\rightarrow\bm{X}\rightarrow\bm{Z}$ forms a Markov chain, which satisfies $I(\bm{Z};\bm{M}|\bm{X})=0$.
-To interpret the second graph, given an MDP $M\sim\bm{M}$, the state-action component of $\bm{X}$ is fully captured by the behavior policy $\pi_{\beta}$: $\bm{s}\sim\mu_{\pi_{\beta}}(\bm{s}),\bm{a}\sim\pi_{\beta}$.
+To interpret the second graph, given an MDP $M\sim\bm{M}$, the state-action component of $\bm{X}$ is primarily captured by the behavior policy $\pi_{\beta}$: $\bm{s}\sim\mu_{\pi_{\beta}}(\bm{s}),\bm{a}\sim\pi_{\beta}$.
+The only exception is when tasks differ in initial state distribution $\rho_{0}$ or transition dynamics $T$, in which case the state variable $\bm{S}$ also depends on $\bm{M}$.
 We therefore define it as the behavior-related component, which should be minimally causally related (dashed lines) to $\bm{M}$ and $\bm{Z}$.
@@
 Theorem 2.3.
-Let $\equiv$ denote equality up to a constant,
-$\underbrace{I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})}_{\text{primary causality}}\quad\leq\quad I(\bm{Z};\bm{M})\quad\leq\underbrace{I(\bm{Z};\bm{X})}_{\text{primary + lesser causality}}$
+Let $\equiv$ denote equality up to a constant, then
+$\underbrace{I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})}_{\textup{primary causality}}\quad\leq\quad I(\bm{Z};\bm{M})\quad\leq\quad I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+I(\bm{Z};\bm{X}_{b})=\underbrace{I(\bm{Z};\bm{X})}_{\textup{primary + lesser causality}}$
 holds up to a constant, where
   1. 1.
+    $\mathcal{L}_{\textup{FOCAL}}\equiv-I(\bm{Z};\bm{X})$.
+  2. 2.
+    $\mathcal{L}_{\textup{CORRO}}\equiv-I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$.
+  3. 3.
+    $\mathcal{L}_{\textup{CSRO}}\geq-\left((1-\lambda)I(\bm{Z};\bm{X})+\lambda I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})\right)$.
+Proof.
+See Appendix B. ∎
+Theorem 2.3 induces several key observations.
+Firstly, the FOCAL and CORRO objectives operate as an upper bound and a lower bound of the relevant information $I(\bm{Z};\bm{M})$ respectively.
+Since one would like to maximize $I(\bm{Z};\bm{M})$ according to Definition 2.1, CORRO, which maximizes the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$, can effectively optimize $I(\bm{Z};\bm{M})$ with theoretical assurance.
+However, FOCAL which maximizes the upper bound $I(\bm{Z};\bm{X}_{t},\bm{X}_{b})$ provides no guarantee for $I(\bm{Z};\bm{M})$.
+Mathematically since $I(\bm{Z};\bm{X})=I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+I(\bm{Z};\bm{X}_{b})$, maximizing the FOCAL objective may instead significantly elevates the lesser causality $I(\bm{Z};\bm{X}_{b})$, which is undesirable since it contains spurious correlation between the task representation $\bm{Z}$ and behavior policy $\pi_{\beta}$.
+This explains why FOCAL is less robust to context shift compared to CORRO.
+Secondly, CSRO as the latest COMRL algorithm, inherently optimizes a linear combination of the FOCAL and CORRO objectives.
+In the $0\leq\lambda\leq 1$ regime, the CSRO objective becomes a convex interpolation of the upper bound $I(\bm{Z};\bm{X})$ and the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ of $I(\bm{Z};\bm{M})$, which in essence, enforces a trade-off between the causal ($\bm{Z}$ with $T\,\&\,\rho_{0}$) and spurious ($\bm{Z}$ with $\pi_{\beta}$) correlation contained in $I(\bm{Z};\bm{X}_{b})$.
+This accounts for the improved performance of CSRO compared to FOCAL and CORRO.
+2.3 Instantiations of UNICORN
+By providing a unified view of pre-existing COMRL algorithms, Theorem 2.3 opens up avenues for novel algorithmic implementations by seeking alternative approximations of the true objective $I(\bm{Z};\bm{M})$.
+To demonstrate the impact of our proposed UNICORN framework, we discuss two instantiations as follows:
 […]
-  2. 2.
+$I(\bm{Z};\bm{M})$ $=H(\bm{M})-H(\bm{M}|\bm{Z})\equiv-H(\bm{M}|\bm{Z})$
+$=\mathbb{E}_{\bm{z}}\mathbb{E}_{M\sim p(M|\bm{z})}\left[\log p(M|\bm{z})\right]=-\mathbb{E}_{\bm{z}}\left[H(\bm{M}|\bm{Z}=\bm{z})\right].$ (7)
+where $H(\cdot)$ is entropy.
+Since in practice, each $\bm{z}^{i}$ of sample $\bm{x}^{i}$ is collected within a specific task $M^{i}$, minimizing the parameterized entropy $H_{\bm{\theta}}(\bm{M}|\bm{Z}=\bm{z}^{i})$ is equivalent to finding an optimal function $p_{\bm{\theta}}(M|\bm{z})$ which correctly assigns the ground-truth label $M^{i}$ to $\bm{z}^{i}$, i.e., optimizing $p_{\bm{\theta}}(M|\bm{z})$ towards a delta function $\delta(M-M^{i})$ for continuous $\bm{M}$ or an indicator function $\mathbbm{1}(M=M^{i})$ for discrete $\bm{M}$.
+This implies that
 […]
-  3. 3.
+Suppose the task label $M$ is given during training and a total of $n_{M}$ training tasks $\{M^{i}\}_{i=1}^{n_{M}}$ are drawn from the task distribution $p(M)$ for meta-training.
+Under this supervised scenario, by substituting Eqn 8 into 7, we have
 […]
+[…]
+which is precisely the negative cross-entropy loss $H(\bm{M},P(\bm{M}|\bm{X}))$ for $n_{M}$-way classification with feature $\bm{z}$ and classifier $p_{\bm{\theta}}$.
+We therefore define the objective of supervised UNICORN as
+$\mathcal{L}_{\textup{UNICORN-SUP}}$ $\coloneqq H(\bm{M},P(\bm{M}|\bm{X}))$
+$=-\mathbb{E}_{\bm{x},\bm{z}\sim q_{\bm{\phi}}(\bm{z}|\bm{x})}\left[\sum_{j=1}^{n_{M}}\mathbbm{1}(M^{i}=M)\log p_{\bm{\theta}}(M^{i}|\bm{z})\right].$ (10)
+Note that $\mathcal{L}_{\textup{UNICORN-SUP}}$ is convex and operates as a finite-sample approximation of $-I(\bm{Z};\bm{M})$, for which we give the following theorem:
+Theorem 2.4 (Concentration bound for supervised UNICORN).
+Denote by $\hat{I}(\bm{Z};\bm{M})$ the empirical estimate of $I(\bm{Z};\bm{M})$ by $n_{M}$ tasks, $\bar{I}(\bm{Z};\bm{M})$ the expectation, then with probability at least $1-\delta$,
+$\left|\hat{I}(\bm{Z};\bm{M})-\bar{I}(\bm{Z};\bm{M})\right|\leq\sqrt{\frac{\textup{Var}(H(\bm{Z}|\bm{M}))}{n_{M}\delta}}.$ (11)
 Proof.
-See Appendix B, Theorem B.1. ∎
-Theorem 3.3 induces several key observations.
-Firstly, the FOCAL and CORRO objective operates as an upper bound and a lower bound of the relevant information $I(\bm{Z};\bm{M})$ respectively.
-Since one would like to maximize $I(\bm{Z};\bm{M})$ according to Definition 3.1, CORRO, which maximizes a lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$, can effectively optimize $I(\bm{Z};\bm{M})$ with theoretical guarantees.
-However, FOCAL which maximizes an upper bound $I(\bm{Z};\bm{X}_{t},\bm{X}_{b})$ is not guaranteed to optimize $I(\bm{Z};\bm{M})$.
-Mathematically since $I(\bm{Z};\bm{X})=I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+I(\bm{Z};\bm{X}_{b})$, maximizing the FOCAL objective may also increase the lesser causality $I(\bm{Z};\bm{X}_{b})$, which is not ideal since it contains spurious correlation between the task representation $\bm{Z}$ and behavior policy $\pi_{\beta}$.
-This explains why FOCAL is less robust to context shift compared to CORRO.
-Secondly, CSRO as a latest COMRL algorithm, inherently optimizes a linear combination of the FOCAL and CORRO objectives.
-In the $0\leq\lambda\leq 1$ regime, the CSRO objective becomes a convex interpolation of the upper bound $I(\bm{Z};\bm{X})$ and the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ of $I(\bm{Z};\bm{M})$, which may form a tighter bound on the real objective $I(\bm{Z};\bm{M})$.
-This explains the improved performance of CSRO compared to FOCAL and CORRO.
-3.3 UNICORN: Implementation of The Information Bottleneck Principle on COMRL
-Given Definition 3.1 of the task representation learning, we now further generalize the pre-existing COMRL algorithms based on the information bottleneck principle (Tishby et al., 2000; Tishby and Zaslavsky, 2015).
-Namely, finding an optimal task representation $\bm{Z}$ can be formulated as the minimization of the following Lagrangian
-$\mathcal{L}_{IB}=I(\bm{Z};\bm{X})-\beta I(\bm{Z};\bm{M})$ (7)
-subject to the Markov chain constraint.
-The positive Lagrange multiplier $\beta$ acts as a tradeoff parameter between the complexity/rate of the representation, $R=I(\bm{Z};\bm{X})$, and the amount of preserved relevant information, $I_{M}=I(\bm{Z};\bm{M})$.
-For implementation, the first term is just FOCAL loss by Theorem 3.3, whereas we propose to approximate the second term as a convex combination of the FOCAL and CORRO loss, analogous to CSRO:
+See Appendix B. ∎
+Figure 3: Meta-learning procedure of UNICORN-SS.
+The supervised variant UNICORN-SUP simply replaces the decoder by a classifier $p_{\bm{\theta}}(M|\bm{z})$ and optimize a cross-entropy loss instead of $\mathcal{L}_{\textup{recon}}$ and $\mathcal{L}_{\textup{FOCAL}}$.
+Theorem 2.4 bounds the finite-sample estimation error of the empirical risk $\hat{I}(\bm{Z};\bm{M})$ with $n_{M}$ task instances drawn from the real task distribution $p(\bm{M})$.
+The supervised UNICORN has the merit of directly estimating and optimizing the real objective $I(\bm{Z};\bm{M})$, which requires explicit knowledge of the task label $M^{i}$ and a substantial amount of task instances according to Theorem 2.4.
+To trade-off computation and performance, we choose to sample 20 training tasks for all RL environments in our experiments.
+Self-supervised UNICORN  In practice, offline RL datasets may often be collected with limited knowledge of the task specifications or labels.
+In this scenario, previous works implement self-supervised learning [31] to obtain effective representation $\bm{Z}$, such as the contrastive-based FOCAL/CORRO to optimize $I(\bm{Z};\bm{X})$/$I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ respectively; or generative approaches like VariBAD [32]/BOReL [13] to reconstruct the trajectories $\bm{X}$ by variational inference, which is equivalent to maximizing $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$.
+By Theorem 2.3, these methods optimize a relatively loose upper/lower bound of $I(\bm{Z};\bm{M})$, which can be improved by a convex combination of the two bounds:
 $I(\bm{Z};\bm{M})\approx\alpha I(\bm{Z};\bm{X})+(1-\alpha)I(\bm{Z};\bm{X}_{t}|\bm{X}_{b}),$ (12)
-Figure 2: Meta-training procedure of UNICORN.
-Built upon the classic COMRL methods FOCAL, UNICORN implements the information bottleneck principle by jointly optimizing the FOCAL loss and a reconstruction loss, using a variational auto-encoder to obtain robust task representation $\bm{z}$.
-where $0\leq\alpha\leq 1$. Substituting Eqn 8 into 7, we may rescale $\mathcal{L}_{IB}$ as
-$\mathcal{L}_{IB}=-\left(I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+\frac{\alpha\beta-1}{(1-\alpha)\beta}I(\bm{Z};\bm{X})\right).$ (9)
-For implementation of $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$, however, it would require additional approximations such as generating negative samples for contrastive learning using a trained model in CORRO, or using a CLUB upper bound of the mutual information in CSRO.
-Instead, we propose to rewrite it according to the chain rule of mutual information (Yeung, 2008):
+where $0\leq\alpha\leq 1$ is a hyperparameter.
+Implementing each term in Eqn 12 allows ample decision choices, such as the contrastive losses in Eqs.
+3, 4 and 6 or autoregressive generation via Decision Transformer [33, 34] or RNN [32].
+For demonstration, in this paper we employ a contrastive objective $\mathcal{L}_{\textup{FOCAL}}$ as in Eq.
+3 for $I(\bm{Z};\bm{X})$ while approximate $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ by reconstruction.
+By the chain rule of mutual information [35]:
 $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ $=I(\bm{X}_{t};\bm{Z},\bm{X}_{b})-I(\bm{X}_{t};\bm{X}_{b})$
@@
 since $I(\bm{X}_{t};\bm{X}_{b})$ is a constant when $\bm{X}_{t}$ and $\bm{X}_{b}$ are drawn from a fixed distribution as in offline RL. Moreover, by definition of mutual information:
-$I(\bm{X_{t}};\bm{Z},\bm{X_{b}})$ $=\int p(\bm{x_{t}},\bm{x_{b}})p(\bm{z}|\bm{x_{t}},\bm{x_{b}})\log\frac{p(\bm{x_{t}}|\bm{z},\bm{x_{b}})}{p(\bm{x_{t}})}$
-$\equiv\int p(\bm{x_{t}},\bm{x_{b}})p(\bm{z}|\bm{x_{t}},\bm{x_{b}})\log p(\bm{x_{t}}|\bm{z},\bm{x_{b}})$
-$\simeq\int_{\bm{x_{t}},\bm{x_{b}},\bm{z}}\underbrace{q_{\bm{\phi}}(\bm{z}|\bm{x_{t}},\bm{x_{b}})}_{\text{encoder}}\underbrace{\log p_{\bm{\theta}}(\bm{x_{t}}|\bm{z},\bm{x_{b}})}_{\text{decoder}}$
-$=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x_{t}},\bm{x_{b}})}\left[\log p_{\bm{\theta}}(\bm{x_{t}}|\bm{z},\bm{x_{b}})\right],$ (11)
-which induces an unbiased estimator of $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ by reconstructing $\bm{X}_{t}$ with a decoder network $p_{\bm{\theta}}(\cdot|\bm{z},\bm{x_{b}})$ conditioning on $\bm{Z}$ and $\bm{X_{b}}$, reminiscent of VAE (Kingma and Welling, 2014).
-We therefore implement the objective of our proposed algorithm UNICORN as a linear combination of this reconstruction loss $\mathcal{L}_{\text{recon}}:=-I(\bm{X_{t}};\bm{Z},\bm{X_{b}})$ and the FOCAL loss $\mathcal{L}_{\text{FOCAL}}=-I(\bm{Z};\bm{X})$:
-$\mathcal{L}_{\text{UNICORN}}$ $:=\mathcal{L}_{IB}=\mathcal{L}_{\text{recon}}+\frac{\alpha\beta-1}{(1-\alpha)\beta}\mathcal{L}_{\text{FOCAL}}$
-$=-\left(I(\bm{X}_{t};\bm{Z},\bm{X}_{b})+\frac{\alpha\beta-1}{(1-\alpha)\beta}I(\bm{Z};\bm{X})\right).$ (12)
-UNICORN can be interpreted as a generalization of the FOCAL and CORRO objecitves by effectively striking a balance between the primary causality $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ and the lesser causality $I(\bm{Z};\bm{X}_{b})$.
-Figure 2 gives an overview of our framework, which enjoys three core advantages:
-  • Robustness and bias-reduction.
-By optimizing a reconstruction loss $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ equivalent to the CORRO objective $I(\bm{Z};\bm{X_{t}}|\bm{X_{b}})$, UNICORN mitigates the spurious correlation in FOCAL, leading to more robust task representation $\bm{Z}$.
-It is also verified in Section 4.5 that UNICORN is robust to a wide range of data qualities.
-Moreover, the unbiased estimator in Eqn 11 circumvents additional approximations as in CORRO and CSRO, making UNICORN a bias-reduced and improved version of previous SOTAs.
-  • Broader applicability domain.
-Unlike FOCAL/CORRO which assumes uniform task distribution, the unbiased estimator in Eqn 11 can be efficiently implemented by Monte Carlo sampling on offline datasets drawn from any distribution.
-Additionally, we will show in Section 5.1 that UNICORN is transferable to other model architectures such as Decision Transformer (DT).
-  • General and unified paradigm.
-Derived from the general information bottleneck principle, our framework encompasses a series of milestones in COMRL including FOCAL, CORRO, CSRO and a concurrent method GENTLE (Zhou et al., 2023), which solely optimizes the reconstruction loss with a similar implementation.
-A unified picture of these pre-existing methods being special cases of UNICORN in the $\alpha$-$\beta$ plane is depicted in Figure 3.
-A hyper-parameter study showing the effect of the UNICORN feasible region is presented in Appendix C.1.
-Moreover, the decoder $p_{\theta}(\bm{x}_{t}|\bm{z},\bm{x}_{b})$ can operate as a model of the MDP to generate additional data using customized latent $\bm{z}$, which is explored in Section 5.2.
-In summary, UNICORN not only provides a unified framework of model-free COMRL, it may also be exploited and lead to novel model-based algorithms.
+$I(\bm{X_{t}};\bm{Z},\bm{X_{b}})$ $=\mathbb{E}_{\bm{x_{t}},\bm{x_{b}},\bm{z}}\left[\log\frac{p(\bm{x_{t}}|\bm{z},\bm{x_{b}})}{p(\bm{x_{t}})}\right]$
+$\equiv\mathbb{E}_{\bm{x_{t}},\bm{x_{b}},\bm{z}}\left[\log p(\bm{x_{t}}|\bm{z},\bm{x_{b}})\right]$
+$\geq\mathbb{E}_{\bm{x_{t},x_{b}},\bm{z}\sim q_{\bm{\phi}}(\bm{z}|\bm{x_{t}},\bm{x_{b}})}\left[\log p_{\bm{\theta}}(\bm{x_{t}}|\bm{z},\bm{x_{b}})\right],$ (14)
+which induces a generative objective $\mathcal{L}_{\textup{recon}}\coloneqq-I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ by reconstructing $\bm{X}_{t}$ with a decoder network $p_{\bm{\theta}}(\cdot|\bm{z},\bm{x_{b}})$ conditioning on $\bm{Z}$ and $\bm{X_{b}}$.
+As a result, the proposed unsupervised UNICORN objective can be rescaled as Eqn 15.
+The influence of the hyper-parameter $\frac{\alpha}{1-\alpha}$ is shown in Appendix C.4.
+$\mathcal{L}_{\textup{UNICORN-SS}}\coloneqq\mathcal{L}_{\textup{recon}}+\frac{\alpha}{1-\alpha}\mathcal{L}_{\textup{FOCAL}}.$ (15)
+We summarize our learning procedure in Fig.
+3 with pseudo-code in Algorithms 1 and 2.
+A holistic comparison of our proposed algorithms with related contextual meta-RL methods is shown in Table 1.
+The extra KL divergence can be interpreted as the result of a variational approximation to an information bottleneck [36, 37] that constrains the mutual information between $\bm{Z}$ and $\bm{X}$, which we found unnecessary in our offline setting (see ablation in Table 6.) Behavior regularized actor critic [38] is employed to tackle the bootstrapping error [39] for downstream offline RL implementation.
+Table 1: Comparison between UNICORN instantiations and related existing contextual meta-RL methods.
+For clarity, "Representation Learning Objective" only lists the loss functions of $\bm{Z}$ that are independent of the downstream RL tasks.
+Note that $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})\equiv I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ holds only for offline RL.
+Method Setting Representation Learning Objective Implementation Context $\bm{X}$ UNICORN-SUP Offline $I(\bm{Z};\bm{M})$ Predictive Transition UNICORN-SS Offline $\alpha I(\bm{Z};\bm{X})+(1-\alpha)I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ Contrastive+Generative Transition FOCAL [11, 21] Offline $I(\bm{Z};\bm{X})$ Contrastive Transition CORRO [23] Offline $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ Contrastive Transition CSRO [22] Offline $(1-\lambda)I(\bm{Z};\bm{X})+\lambda I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ Contrastive Transition GENTLE [40] Offline $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ Generative Transition BOReL [13] Offline $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})-D_{\textup{KL}}(q_{\phi}(\bm{Z}|\bm{X})||p_{\theta}(\bm{Z}))$ Generative Trajectory VariBAD [32] Online $I(\bm{X}_{t};\bm{Z},\bm{X}_{b})-D_{\textup{KL}}(q_{\phi}(\bm{Z}|\bm{X})||p_{\theta}(\bm{Z}))$ Generative Trajectory PEARL [7] Online $-D_{\textup{KL}}(q_{\phi}(\bm{Z}|\bm{X})||p_{\theta}(\bm{Z}))$ N/A Transition ContraBAR [41] Offline&Online $I(\bm{Z};\bm{X}_{t}|\bm{A})$ Contrastive Trajectory
 3 Experiments
-Our main experiments are organized to address three primary inquiries regarding the core advantages of UNICORN: (1) How does UNICORN perform on in-distribution tasks (bias-reduction)?
-(2) How well can UNICORN generalize to data collected by OOD behavior policies (robustness)?
-(3) Can UNICORN outperform consistently across datasets of different qualities (robustness)?
-The performance are measured by the average return across all test tasks.
-Figure 3: The hyper-parameter plane of UNICORN.
-Empirically, we find that UNICORN works properly only when $(\alpha\beta-1)/(1-\alpha)\beta>0$, which further constrains the feasible region of UNICORN (shaded) to be bounded by $\alpha\beta\geq 1$, $\alpha\leq 1$ and $\beta\geq 1$.
-Intriguingly, pre-existing methods FOCAL (red) and CORRO/CSRO/GENTLE (blue) constitute the two boundaries of the feasible region, which can therefore be interpreted as special cases of UNICORN.
-4.1 Environments and Datasets
-We adopt MuJoCo (Todorov et al., 2012) and MetaWorld (Yu et al., 2020a) benchmarks to evaluate our method.
-HalfCheetah-Dir, HalfCheetah-Vel, Ant-Dir are multi-task MuJoCo benchmarks where tasks differ in reward function.
-For HalfCheetah-Dir and Ant-Dir, the agent is rewarded with high velocity along the goal direction.
-For HalfCheetah-Vel, the agent receives an $L_{2}$ penalty with respect to the target velocity.
-Hopper-Param, Walker-Param are multi-task MuJoCo benchmarks where tasks differ in transition dynamics. The physical parameters of body mass, inertia, damping, and friction vary for each task.
-Reach is a multi-task MetaWorld benchmark where tasks differ in reward function. The agent controls the end-effector to reach a target position in 3D space.
-Following the protocol of CORRO, we sample 20 training tasks and 20 testing tasks from the task distribution.
-We train SAC agent (Haarnoja et al., 2018) independently for each task and collect offline datasets of mixed quality using all logged behavior policies.
-4.2 Baselines
-We compare UNICORN with six competitive OMRL algorithms. These baselines are categorized as context-based, gradient-based or transformer-based methods.
-FOCAL, CORRO, CSRO are context-based methods which can be seen as special cases or approximations of UNICORN, see details in Section 3.
-Supervised is a context-based method, directly using actor-critic loss to train the policy/value networks and task encoder end-to-end. We find it a competitive baseline across all benchmarks.
-MACAW (Mitchell et al., 2021) is a gradient-based method, using supervised advantage-weighted regression for both the inner and outer loop of meta-training.
-Prompt-DT (Xu et al., 2022) is a transformer-based method, taking context as the prompt of a Decision Transformer (DT) (Chen et al., 2021; Janner et al., 2021) to solve OMRL as a conditional sequence modeling problem.
-We perform our evaluation in a few-shot manner.
-For gradient-based methods, we first update the meta-policy with a batch of context, then evaluate the agent.
-For transformer-based methods, we utilize $\bm{z}$ inferred from a batch of context as the prompt to condition the auto-regressive rollout.
-4.3 Few-Shot Generalization to In-Distribution Data
-Figure 4: Testing returns of UNICORN against baselines on six benchmarks.
-Solid curves refer to the mean performance of trials over 6 random seeds, and the shaded areas characterize the standard deviation of these trials.
+For brevity, we name our proposed supervised and self-supervised algorithms UNICORN-SUP and UNICORN-SS respectively.
+To evaluate their effectiveness, our main experiments are organized to address three primary inquiries regarding the core advantages of UNICORN: (1) How does UNICORN perform on in-distribution tasks?
+(2) How well can UNICORN generalize to data collected by OOD behavior policies?
+(3) Can UNICORN outperform consistently across datasets of different qualities?
+The performance is measured by the average return across 20 testing tasks randomly sampled for each environment.
+3.1 Experimental Setup
+We adopt MuJoCo [25] and MetaWorld [3] benchmark to evaluate our method, involving six robotic locomotion tasks which require adaptation across reward functions (HalfCheetah-Dir, HalfCheetah-Vel, Ant-Dir, Reach) or across dynamics (Hopper-Param, Walker-Param).
+We compare UNICORN with six competitive OMRL algorithms.
+These baselines are categorized as context-based, gradient-based and transformer-based methods:
 Table 2: Average testing returns of UNICORN against baselines on datasets collected by IID and OOD behavior policies.
@@
 The best is bolded and the second best is underlined.
-Algorithm HalfCheetah-Dir HalfCheetah-Vel Ant-Dir Hopper-Param Walker-Param Reach
-IID OOD IID OOD IID OOD IID OOD IID OOD IID OOD
-UNICORN 1312.59±31.72 1301.17±25.98 -21.88±0.65 -93.57±4.85 266.85±14.54 235.55±17.69 315.90±5.67 304.45±10.88 407.83±75.58 392.48±74.45 2774.67±241.04 2603.68±182.85
-CSRO 1180.48±228.10 458.08±253.00 -27.98±0.90 -102.44±5.17 276.18±19.37 233.33±12.37 309.63±6.50 301.12±9.99 310.47±58.29 278.94±65.37 2719.50±234.55 2800.93±182.06
-CORRO 704.20±449.46 244.95±145.83 -37.11±2.60 -112.29±1.62 147.93±12.85 120.52±12.48 283.04±7.89 271.48±12.79 277.07±37.91 213.43±48.33 2467.83±174.57 2321.87±327.19
-FOCAL 1186.03±271.58 860.58±252.96 -22.22±0.88 -96.72±1.69 217.10±29.14 173.16±24.37 301.84±4.48 296.87±12.76 308.05±98.42 286.34±91.08 2423.67±255.95 2315.56±303.03
-Supervised 961.62±355.84 781.51±429.12 -23.81±0.71 -104.31±0.65 237.52±38.81 202.19±37.56 306.54±9.99 293.65±7.99 256.52±60.39 209.64±27.72 2489.37±248.05 2283.26±205.36
-MACAW 1154.83±10.11 449.56±6.54 -56.04±2.42 -188.44±0.91 26.29±2.80 0.18±0.31 218.52±6.01 205.39±2.47 141.25±8.83 129.63±5.09 2430.67±157.42 1728.01±78.98
-Prompt-DT 1176.54±39.64 -25.10±8.74 -118.14±65.59 -248.64±21.38 1.00±0.38 0.04±0.11 234.19±5.35 202.51±4.96 184.73±8.57 155.78±16.64 2164.86±85.20 1895.51±110.64
+Algorithm HalfCheetah-Dir HalfCheetah-Vel Ant-Dir Hopper-Param Walker-Param Reach IID OOD IID OOD IID OOD IID OOD IID OOD IID OOD UNICORN-SS 1307±26 1296±24 -22±1 -94±5 267±14 236±18 316±6 304±11 419±44 407±46 2775±241 2604±183 UNICORN-SUP 1296±20 1130±76 -25±3 -91±5 250±4 239±16 312±4 302±12 322±28 312±39 2681±111 2641±140 CSRO 1180±228 458±253 -28±1 -102±5 276±19 233±12 310±6 301±10 310±58 279±65 2720±235 2801±182 CORRO 704±450 245±146 -37±3 -112±2 148±13 120±12 283±8 272±13 277±38 213±48 2468±175 2322±327 FOCAL 1186±272 861±253 -22±1 -97±2 217±29 173±24 302±4 297±13 308±98 286±91 2424±256 2316±303 Supervised 962±356 782±429 -24±1 -104±1 238±39 202±38 306±10 294±8 256±60 210±28 2489±248 2283±205 MACAW 1155±10 450±6 -56±2 -188±1 26±3 0±0 218±6 205±2 141±9 130±5 2431±157 1728±79 Prompt-DT 1176±40 -25±9 -118±66 -249±21 1±0 0±0 234±5 202±5 185±9 156±17 2165±85 1896±111
+FOCAL, CORRO, CSRO are context-based methods that can be seen as special cases or approximations of UNICORN, see details in Section 2.
+Supervised is a context-based method, directly using actor-critic loss to train the policy/value networks and task encoder end-to-end. We find it a competitive baseline across all benchmarks.
+MACAW [12] is a gradient-based method, using supervised advantage-weighted regression for both the inner and outer loop of meta-training.
+Prompt-DT [42] is a transformer-based method, taking context as the prompt of a Decision Transformer (DT) [33, 34] to solve OMRL as a conditional sequence modeling problem.
+Figure 4: Testing returns of UNICORN against baselines on six benchmarks.
+Solid curves refer to the mean performance of trials over 6 random seeds, and the shaded areas characterize the standard deviation of these trials.
+We perform our evaluation in a few-shot manner.
+For gradient-based methods, we first update the meta-policy with a batch of testing context and then evaluate the agent.
+For transformer-based methods, we utilize the trajectory segment as the prompt to condition the auto-regressive rollout.
+3.2 Few-Shot Generalization to In-Distribution Data
+Denote by $\{\pi_{\beta}^{i,t}\}_{i=1:N_{M},t=0:N}$ the checkpoints of behavior policies for all tasks logged during data collection (i.e., training SAC agents), where $i$ refers to the task index and $t$ represents training iteration.
+In the in-distribution (IID) setting, for each test task $M^{j}$, we use behavior policies only from the same tasks $\{\pi_{\beta}^{j,t}\}_{j,t=0:N}$ to collect the context and the conditioned rollouts, as visualized in Figure 1.
 Figure 4 illustrates the learning curves of UNICORN vs.
-all baselines.
-IID entries of Table 1 depicts the performance of in-distribution generalization where the context for each task is collected by the behavior policies trained on the same task.
-The results demonstrate that UNICORN almost unanimously achieves SOTA performance across all benchmarks.
-The observation that FOCAL, CORRO, and CSRO achieve asymptotic performance close to UNICORN is also expected, since they are special cases of UNICORN according to our theory.
+all baselines, which correspond to the IID entries of Table 2.
+The results demonstrate that UNICORN variants, especially UNICORN-SS, consistently achieve SOTA performance across all benchmarks.
+The observation that the asymptotic performance of CSRO is comparable to UNICORN is expected since it also optimizes a linear combination of the lower and upper bound of $I(\bm{Z};\bm{M})$ by Theorem 2.3, with a different implementation choice.
 Notably, UNICORN exhibits remarkable stability, especially on HalfCheetah-Dir and Walker-Param, where the other context-based methods suffer a significant performance decline during the later training process.
@@
 3.3 Few-Shot Generalization to Out-of-Distribution Behavior Policies
-Denote by $\{\pi_{\beta}^{i,t}\}_{i,t}$ the checkpoints of behavior policies for all tasks logged during data collection (i.e., training SAC agents), where $i$ refers to the task index and $t$ represents training iteration.
-For each test task $M^{j}$, we use every individual behavior policy in $\{\pi_{\beta}^{i,t}\}_{i,t}$ to collect the OOD contexts and the conditioned rollouts, as visualized in Figure 1.
+To evaluate OOD generalization, for each test task $M^{j}$, we use all behavior policies in $\{\pi_{\beta}^{i,t}\}_{{i=1:N_{M},t=0:N}}$ to collect the OOD contexts and the conditioned rollouts, as visualized in Figure 1.
 The OOD testing performance is measured by averaging returns across all testing tasks, as shown in the OOD entries of Table 2.
 While all methods suffer notable decline when facing OOD context, UNICORN maintains the strongest performance by a large margin.
-Another observation is that the context-based methods are in general significantly more robust compared to the gradient-based MACAW and transformer-based Prompt-DT, which suggests that context-based paradigm should be prioritized for research of generalizable OMRL.
+Another observation is that the context-based methods are in general significantly more robust compared to the gradient-based MACAW and transformer-based Prompt-DT.
 This also delineates the storyline of COMRL development from 2021 to date
@@
 3.4 Influence of Data Quality
+To test whether the UNICORN framework can be applied to different data distributions, we collect three types of datasets whose size is equal to the mixed dataset used in Section 3.2 and 3.3: random, medium and expert, which are characterized by the quality of the behavior policies.
+As shown in Table 3, UNICORN demonstrates SOTA performance on all types of datasets.
+Notably, despite CSRO having better IID results than UNICORN on Ant-Dir under the mixed distribution, UNICORN surpasses CSRO by a large margin under the narrow distribution (medium and expert) and random distribution.
 Table 3: UNICORN vs. baselines on Ant-Dir datasets of various qualities. Each result is averaged by 6 random seeds. The best is bolded and the second best is underlined.
-Algorithm Random Medium Expert
-IID OOD IID OOD IID OOD
-UNICORN 81.31±17.80 62.20±6.15 219.86±23.04 242.74±10.40 278.82±10.46 261.83±13.49
-CSRO 1.55±2.65 -0.20±0.52 165.65±9.89 198.15±17.35 251.77±39.16 201.57±45.29
-CORRO 0.87±0.53 -0.05±0.06 8.16±5.02 -7.01±2.37 -3.90±9.59 -13.59±9.06
-FOCAL 67.00±26.33 43.65±10.08 171.10±84.26 187.06±86.45 229.02±42.07 246.27±20.21
-Supervised 64.97±5.57 46.90±11.86 149.21±50.21 109.85±80.09 249.11±33.47 214.55±59.56
-MACAW 3.412±0.74 -0.27±0.25 27.69±2.17 1.34±0.75 87.86±42.69 0.64±1.49
-Prompt-DT 0.54±0.29 0.02±0.09 2.14±4.36 0.33±0.57 77.94±14.67 1.24±2.18
-To test whether the UNICORN framework can be applied to different data distributions, we collect three types of datasets whose size is equal to the mixed dataset used in Section 4.3 and 4.4: random, medium and expert, which are characterized by the quality of the behavior policies.
-Each dataset is collected by one SAC agent.
-For random datasets, the SAC agent is randomly initialized and untrained.
-As shown in Table 2, UNICORN demonstrates SOTA performance on all types of datasets.
-Notably, despite CSRO having better IID results than UNICORN on Ant-Dir under the mixed data distribution, UNICORN surpasses CSRO by a large margin under the narrow data distribution (medium and expert) and random data distribution.
+Algorithm Random Medium Expert IID OOD IID OOD IID OOD UNICORN-SS 81±18 62±6 220±23 243±10 279±10 262±13 UNICORN-SUP 75±15 60±5 140±11 126±32 247±15 229±19 CSRO 2±3 0±1 166±10 198±17 252±39 202±45 CORRO 1±1 0±0 8±5 -7±2 -4±10 -14±9 FOCAL 67±26 44±10 171±84 187±86 229±42 246±20 Supervised 65±6 47±12 149±50 110±80 249±33 215±60 MACAW 3±1 0±0 28±2 1±1 88±43 1±1 Prompt-DT 1±0 0±0 2±4 0±1 78±15 1±2
 We observe that CORRO fails on all three datasets.
@@
 4 Discussion
-This section presents more empirical evidence on the broader applicability of the UNICORN framework.
+This section presents more empirical evidence on the applicability of the UNICORN framework.
 4.1 Is UNICORN Model-Agnostic?
-Since UNICORN tackles task representation learning from a general information theoretic perspective, it is in principle model-agnostic.
+Table 4: DT implementation of COMRL on HalfCheetah-Dir and Hopper-Param. Each result is averaged by 6 random seeds.
+Algorithm HalfCheetah-Dir Hopper-Param IID OOD IID OOD UNICORN-SS 1307±26 1296±24 316±6 304±11 UNICORN-SS-DT 1233±10 1186±43 304±4 291±4 UNICORN-SUP-DT 1227±21 1065±57 308±6 297±2 FOCAL-DT 1209±33 652±36 293±4 284±5 Prompt-DT 1177±40 -25±9 234±5 203±5
+As UNICORN tackles task representation learning from a general information theoretic perspective, it is in principle model-agnostic.
 Hence a straightforward idea is to transfer UNICORN to other model architectures like DT, which we envision to be a promising backbone for large-scale training of RL foundation models.
 We employ a simple implementation by embedding the task representation vector $\bm{z}$ as the first token in sequence to prompt the learning and generation of DT.
-Instead of the unlearnable prompt used in the original Prompt-DT, we enforce task representation learning of the prompt by optimizing $\mathcal{L}_{\text{UNICORN}}$ and $\mathcal{L}_{\text{FOCAL}}$.
-We name these DT implementations as UNICORN-DT and FOCAL-DT respectively.
+Instead of the unlearnable prompt used in the original Prompt-DT, we enforce task representation learning of the prompt by optimizing $\mathcal{L}_{\text{UNICORN-SS}}$, $\mathcal{L}_{\text{UNICORN-SUP}}$ and $\mathcal{L}_{\text{FOCAL}}$.
+We name these variants as UNICORN-SS-DT, UNICORN-SUP-DT and FOCAL-DT respectively.
 As shown in Table 4 and Figure 7, applying UNICORN and FOCAL on DT results in significant improvement in both IID and OOD generalization, with UNICORN being the superior option.
-Despite the gap between our implementation of UNICORN-DT and its MLP counterpart in terms of asymptotic performance, we expect UNICORN-DT to extrapolate favorably in the regime of greater dataset size and model parameters due to the scaling law of transformer (Kaplan et al., 2020) and DT (Lee et al., 2022).
+Despite the gap between our implementation of UNICORN-DT and its MLP counterpart in terms of asymptotic performance, we expect UNICORN-DT to extrapolate favorably in the regime of greater dataset size and model parameters due to the scaling law of transformer [43] and DT [44].
 We leave this verification to future work.
-Table 3: DT implementation of COMRL on HalfCheetah-Dir and Hopper-Param. Each result is averaged by 6 random seeds.
-Algorithm HalfCheetah-Dir Hopper-Param
-IID OOD IID OOD
-UNICORN 1312.59±31.72 1301.17±25.98 315.90±5.67 304.45±10.88
-UNICORN-DT 1233.19±10.42 1185.76±42.71 303.66±3.51 290.88±3.68
-FOCAL-DT 1208.58±33.38 652.14±35.86 293.25±4.17 283.86±4.81
-Prompt-DT 1176.54±39.64 -25.10±8.74 234.19±5.35 202.51±4.96
 4.2 Can UNICORN be Exploited for Model-Based Paradigms?
-As alluded in Section 3.3, UNICORN can potentially enable a model-based paradigm by generating data using the decoder $p_{\theta}(\bm{x}_{t}|\bm{z},\bm{x}_{b})$ and customized latent $\bm{z}$, which can be especially useful for generalization to OOD tasks.
+Figure 5: Testing returns for OOD tasks. The learning curve is averaged by 6 random seeds.
+With decoder $p_{\theta}(\bm{x}_{t}|\bm{z},\bm{x}_{b})$, UNICORN-SS can potentially enable a model-based paradigm by generating data using customized latent $\bm{z}$, which can be especially useful for generalization to OOD tasks.
 We implement this idea on Ant-Dir by sorting the tasks according to the polar angle of their goal direction ($\theta\in[0,2\pi)$), and construct task-level OOD by taking 20 consecutive tasks in the middle for training and the rest for testing (Figure 1).
@@
 The imaginary rollouts $\{(\bm{s}_{t},\bm{a}_{t},\bm{s}^{\prime}_{t},r_{t})\}_{t=1}^{n}$ are used during the training of RL agent.
-We adopt the conventional ensemble technique in model-based RL (Yu et al., 2020b; Zhang et al., 2023) to stabilize the training process.
-As shown in Figure 5, UNICORN-enabled model-based RL is the only method to achieve positive performance in this extremely challenging scenario where context shift is induced by OOD tasks.
-Figure 5: Testing returns for OOD tasks. The learning curve is averaged by 6 random seeds.
-6 Conclusion
-In this paper, we unify three major context-based offline meta-RL algorithms, FOCAL, CORRO and CSRO (potentially a lot more) into a single framework based on the information bottleneck principle.
+We adopt the conventional ensemble technique in model-based RL  [45, 46] to stabilize the training process.
+As shown in Figure 5, UNICORN-SS-enabled model-based RL is the only method to achieve positive performance in this extremely challenging scenario where context shift is induced by OOD tasks.
+5 Conclusion & Limitation
+In this paper, we unify three major context-based offline meta-RL algorithms, FOCAL, CORRO and CSRO (potentially a lot more) into a single framework from an information theoretic representation learning perspective.
 Our theory offers valuable insight on how these methods are inherently connected and incrementally evolved in pursuit of better generalization against context shift.
-Based on the proposed framework, we arrive at a novel algorithm called UNICORN which is demonstrated to be remarkably robust and broadly applicable through extensive experiments.
-We believe our study offers potential for new implementations, optimality bounds and algorithms for both fully offline and offline-to-online COMRL paradigms.
-We leave these directions for future work.
+Based on the proposed framework, we instantiate two novel algorithms called UNICORN-SUP and UNICORN-SS which are demonstrated to be remarkably robust and broadly applicable through extensive experiments.
+We believe our study offers potential for new implementations, optimality bounds and algorithms for both fully-offline and offline-to-online COMRL paradigms.
+Mathematically, we expect our framework to be able to incorporate almost all existing COMRL methods that focus on task representation learning.
+Since according to Definition 2.1, as long as the method tries to solve COMRL by learning a sufficient statistic of w.r.t , it will eventually come down to optimizing an information-theoretic objective equivalent to $I(\bm{Z},\bm{M})$, or a lower/upper bound like the ones introduced by Theorem 2.3, up to some regularizations or constraints.
+Limitation Since our framework assumes a decoupling of task representation learning and offline policy optimization, it does not directly elucidate how high-quality representations are indicative of higher downstream RL performance.
+Another limitation would be the scale of the experiments.
+Our offline datasets cover at most 40 tasks and $\sim 450k$ transitions, which might be the reason why UNICORN-SUP is inferior to UNICORN-SS ($n_{M}$ too small in Theorem 2.4).
+However, we expect our conclusion to extrapolate to larger task sets, dataset sizes and model parameters, for which the DT variants may demonstrate better scaling.
+Lastly, in principle, the information theoretic formalization in this work should be directly applicable to online RL.
+However, many of our key derivations rely on the static assumption of $\bm{M}$ and $\bm{X}$ (see Appendix B), which are apparently violated in the online scenario.
+Extending our framework to online RL is interesting and nontrivial, which we leave for future work.
+Acknowledgments and Disclosure of Funding
+The authors would like to thank the anonymous reviewers for their insightful comments and suggestions.
+The work described in this paper was supported partially by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project Reference Number: T45-401/22-N), by a grant from the Hong Kong Innovation and Technology Fund (Project Reference Number: MHP/086/21) and by the National Key Research and Development Program of China (No.
+2020YFA0711402).
 Appendix A Pseudo-Code
 Algorithm 1 UNICORN Meta-training
-0:  Offline Datasets $\bm{X}=\{X^{i}\}_{i=1}^{N_{train}}$, training tasks $\bm{M}=\{M^{i}\}_{i=1}^{N_{train}}$, initialized learned policy $\pi_{\omega}$, Q function $Q_{\psi}$, context encoder $q_{\phi}$, decoder $p_{\theta}$, hyper-parameter $\alpha$, $\beta$, task batch size $N_{tbatch}$
-1:  while not done do
-2:     Sample task batch $\{M^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{M}$ and the corresponding replay buffer =$\mathcal{R}=\{X^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{X}$
-3:     for step in each iter do
-4:        // Train the context encoder and decoder
-5:        Sample context $\{\bm{c}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
-6:        Obtain task embeddings $\{\bm{z}^{j}\sim q_{\phi}(\bm{z}|\bm{c}^{j})\}_{j=1}^{N_{tbatch}}$
-7:        Compute $\mathcal{L}_{\text{UNICORN}}=\mathcal{L}_{\text{recon}}+\frac{\alpha\beta-1}{(1-\alpha)\beta}\mathcal{L}_{\text{FOCAL}}$
-8:        Update $\phi,\theta$ to minimize $\mathcal{L}_{\text{UNICORN}}$
-9:        // Train the actor and critic
-10:        Detach the task embeddings $\{\bm{z}^{j}\}_{j=1}^{N_{tbatch}}$
-11:        Sample training data $\{\bm{d}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
-12:        Compute $\mathcal{L}_{\text{actor}},\mathcal{L}_{\text{critic}}$
-13:        Update $\omega,\psi$ to minimize $\mathcal{L}_{\text{actor}},\mathcal{L}_{\text{critic}}$
-14:     end for
-15:  end while
+1:Offline Datasets $\bm{X}=\{X^{i}\}_{i=1}^{N_{\textup{{train}}}}$, training tasks $\bm{M}=\{M^{i}\}_{i=1}^{N_{\textup{train}}}$, initialized learned policy $\pi_{\bm{\omega}}$, Q function $Q_{\bm{\psi}}$, context encoder $q_{\bm{\phi}}$, classifier/decoder $p_{\bm{\theta}}$, hyper-parameter $\alpha$, task batch size $N_{tb}$, learning rates $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$
+2:while not done do
+3:     Sample task batch $\{M^{j}\}_{j=1}^{N_{tb}}\sim\bm{M}$ and the corresponding replay buffer =$\mathcal{R}=\{X^{j}\}_{j=1}^{N_{tb}}\sim\bm{X}$
+4:     for step in each iter do
+5:         // Train the context encoder and decoder
+6:         Sample context $\{\bm{c}^{j}\}_{j=1}^{N_{tb}}\sim\mathcal{R}$
+7:         Obtain task embeddings $\{\bm{z}^{j}\sim q_{\bm{\phi}}(\bm{z}|\bm{c}^{j})\}_{j=1}^{N_{tb}}$
+8:         Estimate $\mathcal{L}_{\text{UNICORN}}$ ($\mathcal{L}_{\text{UNICORN-SUP}}$ or $\mathcal{L}_{\text{UNICORN-SS}}$)
+9:         $\bm{\phi}\leftarrow\bm{\phi}-\beta_{1}\nabla_{\bm{\phi}}\mathcal{L}_{\text{UNICORN}}$
+10:         $\bm{\theta}\leftarrow\bm{\theta}-\beta_{2}\nabla_{\bm{\theta}}\mathcal{L}_{\text{UNICORN}}$
+11:         // Train the actor and critic
+12:         Detach the task embeddings $\{\bm{z}^{j}\}_{j=1}^{N_{tb}}$
+13:         Sample training data $\{\bm{d}^{j}\}_{j=1}^{N_{tb}}\sim\mathcal{R}$
+14:         Estimate $\mathcal{L}_{\text{actor}},\mathcal{L}_{\text{critic}}$
+15:         $\bm{\omega}\leftarrow\bm{\omega}-\beta_{3}\nabla_{\bm{\omega}}\mathcal{L}_{\text{actor}}$
+16:         $\bm{\psi}\leftarrow\bm{\psi}-\beta_{4}\nabla_{\bm{\psi}}\mathcal{L}_{\text{critic}}$
+17:     end for
+18:end while
 Algorithm 2 UNICORN Meta-testing
-0:  Offline Datasets $\bm{X}=\{X^{i}\}_{i=1}^{N_{test}}$, testing tasks $\bm{M}=\{M^{i}\}_{i=1}^{N_{test}}$, learned policy $\pi_{\omega}$, context encoder $q_{\phi}$
-1:  for each $M^{i}$ do
-2:     Sample context $\bm{c}^{i}\sim X^{i}$
-3:     Obtain task embedding $\bm{z}^{i}\sim q_{\phi}(\bm{z}|\bm{c}^{i})$
-4:     Rollout policy $\pi_{\omega}(\bm{a}|\bm{s},\bm{z}^{i})$ for evaluation
-5:  end for
+1:Offline Datasets $\bm{X}=\{X^{i}\}_{i=1}^{N_{\textup{test}}}$, testing tasks $\bm{M}=\{M^{i}\}_{i=1}^{N_{\textup{test}}}$, trained policy $\pi_{\bm{\omega}}$ and context encoder $q_{\bm{\phi}}$
+2:for each $M^{i}$ do
+3:     Sample context $\bm{c}^{i}\sim X^{i}$
+4:     Obtain task embedding $\bm{z}^{i}\sim q_{\bm{\phi}}(\bm{z}|\bm{c}^{i})$
+5:     Rollout policy $\pi_{\bm{\omega}}(\bm{a}|\bm{s},\bm{z}^{i})$ for evaluation
+6:end for
 Appendix B Proofs
-Theorem B.1.
-Let $\equiv$ denote equality up to a constant,
-$\underbrace{I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})}_{\text{primary causality}}\quad\leq\quad I(\bm{Z};\bm{M})\quad\leq\underbrace{I(\bm{Z};\bm{X})}_{\text{primary + lesser causality}}$
-holds up to a constant, where
-  1. 1.
-[…]
-  2. 2.
-[…]
-  3. 3.
-[…]
+B.1 Proof of Theorem 2.3
+We first prove the following lemma:
+Lemma B.1.
+For COMRL, $I(\bm{Z};\bm{M})\geq I(\bm{Z};\bm{M}|\bm{X}_{b})$; or equivalently, $I(\bm{Z};\bm{M};\bm{X}_{b})\geq 0$, both up to a constant.
 Proof.
-Assuming 2, $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})\equiv-\mathcal{L}_{\text{CORRO}}\leq I(\bm{Z};\bm{M})$ is shown in the Theorem 4.1.
-of CORRO [Yuan and Lu, 2022] and GENTLE [Zhou et al., 2023] with strong assumptions.
-We hereby present a proof with minimal assumptions.
-Given $I(\bm{Z};\bm{M};\bm{X}_{b})\geq 0$ in Assumption 3.2, we have
+$I(\bm{Z};\bm{M}|\bm{X}_{t},\bm{X}_{b})=0$
+$\Longrightarrow$ $I(\bm{Z};\bm{M})-I(\bm{Z};\bm{M}|\bm{X}_{t},\bm{X}_{b})=I(\bm{M};\bm{X}_{b},\bm{X}_{t})-I(\bm{M};\bm{X}_{b},\bm{X}_{t}|\bm{Z})\geq 0$
+$\Longrightarrow$ $I(\bm{M};\bm{X}_{b})+I(\bm{M};\bm{X}_{t}|\bm{X}_{b})-I(\bm{M};\bm{X}_{b}|\bm{Z})-I(\bm{M};\bm{X}_{t}|\bm{X}_{b},\bm{Z})\geq 0$
+$\Longrightarrow$ $I(\bm{M};\bm{X}_{b})-I(\bm{M};\bm{X}_{b}|\bm{Z})\geq\underbrace{-I(\bm{M};\bm{X}_{t}|\bm{X}_{b})}_{\text{const}}+\underbrace{I(\bm{M};\bm{X}_{t}|\bm{X}_{b},\bm{Z})}_{\geq 0}\geq\text{const}.$
+Now, since $I(\bm{Z};\bm{M})-I(\bm{Z};\bm{M}|\bm{X}_{b})=I(\bm{M};\bm{X}_{b})-I(\bm{M};\bm{X}_{b}|\bm{Z})$, we have $I(\bm{Z};\bm{M})-I(\bm{Z};\bm{M}|\bm{X}_{b})=I(\bm{Z};\bm{M};\bm{X}_{b})\geq\text{const}\equiv 0$.
+∎
+With Lemma B.1, we proceed to prove Theorem 2.3 as follows:
+Proof.
+$I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})\equiv-\mathcal{L}_{\text{CORRO}}\leq I(\bm{Z};\bm{M})$ is shown in the Theorem 4.1.
+of CORRO [23] and GENTLE [40] with strong assumptions.
+We hereby present a proof with no assumptions.
+Given $I(\bm{Z};\bm{M};\bm{X}_{b})\geq 0$ (Lemma B.1), we have
 $I(\bm{Z};\bm{M})$ $=I(\bm{Z};\bm{M}|\bm{X}_{b})+I(\bm{Z};\bm{M};\bm{X}_{b})$
@@
 as desired.
-The third and fourth lines are obtained by direct application of the chain rule of mutual information [Yeung, 2008].
+The third and fourth lines are obtained by direct application of the chain rule of mutual information [35].
 All mutual information terms without $\bm{Z}$ are held constant in the fully offline scenario.
-$I(\bm{Z};\bm{M})\leq I(\bm{Z};\bm{X})$ is a direct consequence of the Data Processing Inequality [Yeung, 2008] and the Markov chain $\bm{M}\rightarrow\bm{X}\rightarrow\bm{Z}$ in Assumption 3.2.
+$I(\bm{Z};\bm{M})\leq I(\bm{Z};\bm{X})$ is a direct consequence of the Data Processing Inequality [35] and the Markov chain $\bm{M}\rightarrow\bm{X}\rightarrow\bm{Z}$ in Definition 2.2.
 We now prove the three claims regarding FOCAL, CORRO and CSRO:
@@
     $=\mathbb{E}_{\bm{x},\bm{z}}\left[\log\left(\frac{h(\bm{x},\bm{z})}{\sum_{M^{i}\in\mathcal{M}}h(\bm{x}^{i},\bm{z})}\right)\right]+\log(|\mathcal{M}|)$
-    The first term on RHS is precisely the supervised contrastive learning objective introduced by a variant of FOCAL [Li et al., 2021a], which is equivalent to the negative distance metric learning loss $\mathcal{L}_{\text{FOCAL}}$ with the effect of pushing away embeddings of different tasks while pulling together those from the same task.
+    The first term on RHS is precisely the supervised contrastive learning objective introduced by a variant of FOCAL [21], which is equivalent to the negative distance metric learning loss $\mathcal{L}_{\text{FOCAL}}$ with the effect of pushing away embeddings of different tasks while pulling together those from the same task.
 Therefore we have
@@
     $\mathcal{L}_{\text{CSRO}}\geq(\lambda-1)I(\bm{Z};\bm{X})-\lambda I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$.
-    The CLUB loss in Eqn 6 operates as a upper bound of $I(\bm{Z};\bm{X}_{b})$ [Cheng et al., 2020]. By Eqn 5 we have
+    The CLUB loss in Eqn 6 operates as a upper bound of $I(\bm{Z};\bm{X}_{b})$ [30]. By Eqn 5 we have
     $\mathcal{L}_{\text{CSRO}}$ $=\mathcal{L}_{\text{FOCAL}}+\lambda L_{\text{CLUB}}$
@@
     $=(\lambda-1)I(\bm{Z};\bm{X})-\lambda I(\bm{Z};\bm{X}_{t}|\bm{X}_{b}).$
+B.2 Proof of Theorem 2.4
+Proof.
+Denote by $n_{Z}\coloneqq\sum_{i=1}^{n_{M}}|M^{i}||X^{i}|$ the number of context samples, we have
+$I(\bm{Z};\bm{M})$ $=H(\bm{Z})-H(\bm{Z}|\bm{M})$ (16)
+$\simeq-\sum_{i=1}^{n_{Z}}\log p(\bm{z}_{i})-\mathbb{E}_{M}\left[H(\bm{Z}|\bm{M}=M)\right]$ (17)
+$\simeq-\sum_{i=1}^{n_{Z}}\log p(\bm{z}_{i})-\sum_{j=1}^{n_{M}}H(\bm{Z}|\bm{M}=M^{j}).$ (18)
+Since $n_{Z}\gg n_{M}$, the concentration characteristic of $I(\bm{Z};\bm{M})$ is dominated by the second term.
+If we ignore the approximation error of the first term, by the Chebyshev’s inequality, for any $\epsilon>0$,
+$\textup{Pr}\left(\left|\hat{I}(\bm{Z};\bm{M})-\bar{I}(\bm{Z};\bm{M})\right|\geq\epsilon\right)\leq\frac{\textup{Var}(H(\bm{Z}|\bm{M}))}{n_{M}\epsilon^{2}}.$ (19)
+Or equivalently, with probability at least $1-\delta$,
+$\left|\hat{I}(\bm{Z};\bm{M})-\bar{I}(\bm{Z};\bm{M})\right|\leq\sqrt{\frac{\textup{Var}(H(\bm{Z}|\bm{M}))}{n_{M}\delta}}.$ (20)
+Remark: Consider a practical implementation of $p(\bm{z}|M)$ as a multivariate Gaussian $\mathcal{N}(\bm{\mu}_{M},\bm{\Sigma}_{M})$, which gives $H(\bm{Z}|\bm{M})=\frac{1}{2}\log[(2\pi e)^{\dim Z}|\bm{\Sigma}_{M}|]$ [35].
+Substituting into Eqn 20, we have with probability at least $1-\delta$,
+$\left|\hat{I}(\bm{Z};\bm{M})-\bar{I}(\bm{Z};\bm{M})\right|\leq\sqrt{\frac{\textup{Var}(\log|\bm{\Sigma}_{M}|)}{2n_{M}\delta}}.$ (21)
 Appendix C Further Experiments
-Figure 6: Hyper-parameter study results in all environments, each result is averaged over 6 random seeds.
-C.1 Hyper-parameter Study of $\frac{\alpha\beta-1}{(1-\alpha)\beta}$
-We experiment with $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ from 0 to 10, as shown in Figure 6.
-We observe that UNICORN has an optimal $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ in all environments.
-This observation suggests that it is insufficient to use either FOCAL loss or reconstruction loss alone, emphasizing the necessity of their combination proposed in Eqn 7 and 8.
-Intuitively, for tasks differ in transition dynamics, the lesser causality $I(\bm{Z};\bm{X}_{b})$ should be higher since the state distribution $\mu_{\pi}(\bm{s})$ vary across tasks even with the same behavior policy $\pi$.
-This indicates that the optimal proportion of FOCAL loss in Eqn 12 should be greater on these environments, as it maximizes $I(\bm{Z};\bm{X})$ along with the primary causality $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$.
-Such thought experiment is verified in this study.
-We record the optimal $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ of all benchmarks in Table 5, demonstrating that for tasks differ in reward function (Ant-Dir, HalfCheetah-Dir) and transition dynamics (Hopper-Param, Walker-Param), the optimal value of $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ is around 0.1 and 1 respectively.
-The fact that the optimal $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ does not vanish for Ant-Dir and HalfCheetah-Dir suggests that the lesser causality $I(\bm{Z};\bm{X}_{b})$ should be nonzero even when all tasks share the same transition dynamics.
-We conjecture this is partially due to the fact that the behavior policy we use for each task is also trained on that task, which leads to task-dependent state distribution $\mu_{\pi}(\bm{s})$.
-However, this applies to the IID setting only.
-C.2 UNICORN-0: A Label-free Version
-When $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ equals zero, UNICORN task representation learning reduces to the minimization of the reconstruction loss only, which becomes a label-free algorithm (i.e., it does not require the knowledge of task identities/labels to optimize, as opposed to the contrastive learning in previous methods).
-We name this special case UNICORN-0, which is equivalent to a concurrent method GENTLE [Zhou et al., 2023].
-For a fair comparison, we use a variant of a Bayesian OMRL method BOReL [Dorfman et al., 2021] that does not utilize oracle reward functions, as a label-free baseline.
+C.1 UNICORN-SS-0: A Label-free Version
+When $\alpha=0$, UNICORN-SS task representation learning reduces to the minimization of the reconstruction loss only, which becomes a label-free algorithm (i.e., it does not require the knowledge of task identities/labels to optimize, as opposed to the contrastive learning in previous methods).
+We name this special case UNICORN-SS-0, which is equivalent to a concurrent method GENTLE  [40].
+For a fair comparison, we use a variant of a Bayesian OMRL method BOReL [13] that does not utilize oracle reward functions, as a label-free baseline.
 At test time, we modify BOReL to use offline datasets rather than the online data collected by interacting with the environment to infer the task information [^2] Experiment shows that the usage of offline dataset rather than online data has little effect on the results of BOReL..
-As shown in Table 4, UNICORN-0 is also competitive with BOReL.
-Table 4: UNICORN-0 compared to another label-free COMRL method, BOReL, on Ant-Dir.
+As shown in Table 5, UNICORN-SS-0 is also competitive with BOReL.
+Table 5: UNICORN-SS-0 compared to another label-free COMRL method, BOReL, on Ant-Dir.
 Algorithm Ant-Dir
 IID OOD
-UNICORN-0 219.88±16.24 199.86±9.26
-BOReL 206.4±17.93 186.55±9.54
-C.3 Visualization of Task Representations
-For better interpretation, we visualize the task representations by 2D projection of the embedding vectors via t-SNE [Van der Maaten and Hinton, 2008].
+UNICORN-SS-0 220±16 200±9
+BOReL 206±18 187±10
+Table 6: UNICORN-SS-0 compared to other KL constraint weight variants.
+Algorithm Ant-Dir
+IID OOD
+UNICORN-SS-0 220±16 200±9
+KL-0.1 215±11 189±12
+KL-0.5 202±17 182±13
+KL-1.0 194±11 161±7
+KL-5.0 162±14 143±10
+To further validate that the KL constraint $D_{KL}(q_{\theta}(z|x)|\mathcal{N}(0,I))$ is unnecessary, we weight the KL constraint to UNICORN-SS-0.
+Table 6 shows that KL constraint would reduce the performance under the offline setting.
+C.2 Visualization of Task Embeddings
+For better interpretation, we visualize the task representations by 2D projection of the embedding vectors via t-SNE [47].
 For each test task, we generate a trajectory using each policy in $\{\pi_{\beta}^{i,t}\}_{i,t}$ and infer their task representations $\bm{z}$.
 This setting aligns with our OOD experiments in the main text.
-Since the UNICORN objective operates as a convex combination of the reconstruction loss (UNICORN-0) and the FOCAL loss, we compare representations of UNICORN to these two extremes.
-As shown in Figure 7, compared to UNICORN-0, UNICORN provides evidently more distinguishable task representations for OOD data, emphasizing the necessity of joint optimization with the FOCAL loss.
-On the other hand, despite the seemingly higher quality of task representations of FOCAL, the performance of FOCAL is much worse than UNICORN.
-We speculate that since FOCAL blindly separates embeddings from different tasks, it may fail to capture shared structure between similar tasks, whereas UNICORN is able to do so by generative modeling via the reconstruction loss.
+Since the UNICORN-SS objective operates as a convex combination of the reconstruction loss (UNICORN-SS-0) and the FOCAL loss, we compare representations of UNICORN-SS to these two extremes.
+As shown in Figure 6, compared to UNICORN-SS-0, UNICORN-SS provides evidently more distinguishable task representations for OOD data, emphasizing the necessity of joint optimization with the FOCAL loss.
+On the other hand, despite the seemingly higher quality of task representations of FOCAL, the performance of FOCAL is much worse than UNICORN-SS.
+We speculate that since FOCAL blindly separates embeddings from different tasks, it may fail to capture shared structure between similar tasks, whereas UNICORN-SS is able to do so by generative modeling via the reconstruction loss.
 Figure 6: The 2D projection of the learned task representation space in Ant-Dir.
@@
 Tasks of given goals from 0 to 6 are mapped to rainbow colors, ranging from purple to red.
-C.4 More on UNICORN-DT
+C.3 More on UNICORN-DT
 We provide the learning curves of the results in Section 4.1 here.
-As shown in Figure 8, UNICORN-DT demonstrates both the higher convergence speed and the higher asymptotic performance compared to the vanilla Prompt-DT.
-Figure 8: Test returns of UNICORN-DT against UNICORN and Prompt-DT on 2 benchmarks. The learning curve is averaged by 6 random seeds.
+As shown in Figure 7, UNICORN-SS-DT and UNICORN-SUP-DT demonstrate a faster convergence speed compared to vanilla UNICORN-SS and a higher asymptotic performance compared to vanilla Prompt-DT.
+Figure 7: Test returns of UNICORN-SUP-DT and UNICORN-SS-DT against UNICORN-SS, Prompt-DT and FOCAL-DT on 2 benchmarks. The learning curve is averaged by 6 random seeds.
+C.4 Ablation Study
+Figure 8: Different hyper-parameter settings of $\frac{\alpha}{1-\alpha}$ on Ant-Dir. The learning curve is averaged by 6 random seeds.
+To illustrate the effect of hyper-parameter $\frac{\alpha}{1-\alpha}$ on the asymptotic performance, we set the following ablation study.
+Figure 8 shows that as $\frac{\alpha}{1-\alpha}$ increases, the performance gradually increases but when it goes excessively the performance would decrease (FOCAL means $\frac{\alpha}{1-\alpha}\rightarrow\infty$), which validates our proposed theory.
+We also find that UNICORN-SS can maintain relatively stable performance over a range of $\frac{\alpha}{1-\alpha}$ values, thus alleviating the exhaustion from parameter-tuning.
 Appendix D Experimental Details
@@
 Table 7: Configurations and hyper-parameters used in the training process.
-Configurations Ant-Dir HalfCheetah-Dir HalfCheetah-Vel Hopper-Param Walker-Param Reach
-dataset size 1e5 2e5 2e5 3e5 4.5e5 1e5
-task representation dimension 5 5 5 40 32 2
-optimal $\frac{\alpha\beta-1}{(1-\alpha)\beta}$ 0.15 0.01 0.15 1.5 0.7 0.3
-training steps 200k
-task batch size 16
-RL batch size 256
-context training size 200 500
-learning rate 3e-4
-RL network width 256
-RL network depth 3
-encoder width 64 128 64
+Configurations Ant-Dir HalfCheetah-Dir HalfCheetah-Vel Hopper-Param Walker-Param Reach dataset size 1e5 2e5 2e5 3e5 4.5e5 1e5 task representation dimension 5 5 5 40 32 2 weight $\frac{\alpha}{1-\alpha}$ 0.15 0.15 0.15 1.5 1.5 0.3 training steps 200k task batch size 16 RL batch size 256 context training size 1 trajectory (200 steps) 1 trajectory (500 steps) learning rate 3e-4 RL network width 256 RL network depth 3 encoder width 64 128 64
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: Yes, the main contributions are clearly laid out at the end of introduction. See Section 2 for our theory and Section 3 for our experimental verification.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Yes, see Section 5.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: Yes, see for example Definition 2.2 and Appendix B for a complete proof.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Yes, see Appendix D for experimental details. Source code is provided in the Supplementary Material.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: Yes, see Supplementary Material for the source code and README instructions.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Yes, see for example Section C.4 for hyperparameter study.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: Yes, all figures and tables contain error bars.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: See README in Supplementary Material for computing environment information.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We make sure to preserve anonymity and conform to the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: The societal/general impacts of offline meta-RL are discussed in Section 1.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We only use open-source benchmarks for experiments: MuJoCo (Apache-2.0 license ) and Meta-World (MIT license).
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: The paper does not release new assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subject.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@
 As demonstrations, we propose a supervised and a self-supervised implementation of $I(\bm{Z};\bm{M})$, and empirically show that the corresponding optimization algorithms exhibit remarkable generalization across a broad spectrum of RL benchmarks, context shift scenarios, data qualities and deep learning architectures.
-This work lays the information theoretic foundation for COMRL methods, leading to a better understanding of task representation learning in the context of reinforcement learning  [^1] The open-source code is at https://github.com/betray12138/UNICORN.git..
+This work lays the information theoretic foundation for COMRL methods, leading to a better understanding of task representation learning in the context of reinforcement learning.
+Given its generality, we envision our framework as a promising offline pre-training paradigm of foundation models for decision making.
+[^1] Source code: https://github.com/betray12138/UNICORN.git.
 1 Introduction
 The ability to swiftly learn and generalize to new tasks is a hallmark of human intelligence.
-In pursuit of this high level of artificial intelligence (AI), the paradigm of meta-reinforcement learning (RL) proposes to train AI agents in a trial-and-error manner by interacting with multiple external environments.
+In pursuit of this high-level artificial intelligence (AI), the paradigm of meta-reinforcement learning (RL) proposes to train AI agents in a trial-and-error manner by interacting with multiple external environments.
 In order to quickly adapt to the unknown, the agents need to integrate prior knowledge with minimal experience (namely the context) collected from the new tasks or environments, without over-fitting to the new data.
 This meta-RL mechanism has been adopted in many applications such as games [1, 2], robotics [3, 4] and drug discovery [5].
-However, for data collection, classical meta-RL usually requires enormous online explorations of the environments [6, 7], which is impractical in many safety-critical scenarios such as healthcare [8] and robotic manipulation [9].
-As a remedy, offline RL [10] enables agents to learn from logged experience only, therefore circumventing risky or costly online interactions.
-Recently, offline meta-RL (OMRL) [11, 12, 13] has emerged as a novel paradigm to significantly extend the applicable range of RL by "killing two birds in one stone": it builds powerful agents that can quickly learn and adapt by meta-learning, while leveraging offline RL mechanism to ensure a safe and efficient optimization procedure.
+However, for data collection, classical meta-RL usually requires enormous online explorations of the environments [6, 7], which is impractical in many safety-critical scenarios such as healthcare [8], autonomous driving [9] and robotic manipulation [10].
+As a remedy, offline RL [11] enables agents to learn from logged experience only, thereby circumventing risky or costly online interactions.
+Recently, offline meta-RL (OMRL) [12, 13, 14] has emerged as a novel paradigm to significantly extend the applicability of RL by "killing two birds in one stone": it builds powerful agents that can quickly learn and adapt by meta-learning, while leveraging offline RL mechanism to ensure a secure and efficient optimization procedure.
 In the context of classical supervised or self-supervised learning, which is de facto offline, OMRL is reminiscent of the multi-task learning [15], meta-training [6] and fine-tuning [16, 17, 18] of pre-trained large models.
@@
 To alleviate the problem, follow-up works such as CORRO [24] reformulates the task representation learning of COMRL as maximizing the mutual information $I(\bm{Z};\bm{M})$ between the task variable $\bm{M}$ and its latent $\bm{Z}$.
-It then approximates $I(\bm{Z};\bm{M})$ by an InfoNCE [24] contrastive loss, where the positive and negative pairs are conditioned on the state-action tuples $(\bm{s},\bm{a})$.
+It then approximates $I(\bm{Z};\bm{M})$ by an InfoNCE [25] contrastive loss, where the positive and negative pairs are conditioned on the same state-action tuples $(\bm{s},\bm{a})$.
 Inspired by CORRO, a recently proposed method CSRO [23] introduces an additional mutual information term between $\bm{Z}$ and $(\bm{s},\bm{a})$.
@@
 [14] concurrently identified a related problem which they termed MDP ambiguity in the context of Bayesian offline RL.
-An example is illustrated in Figure 1.
+An example is illustrated in Fig.
+1.
 The problem is further exacerbated in the fully offline setting, as the testing distribution is fixed and cannot be augmented by online exploration, allowing no theoretical guarantee for the context shift.
@@
 We start with a formal definition of task representation learning in COMRL:
-Definition 2.1.
-Given an input context variable $\bm{X}\in\mathcal{X}$ and its associated task/MDP random variable $\bm{M}\in\mathcal{M}$, task representation learning in COMRL aims to find a sufficient statistics $\bm{Z}$ of $\bm{X}$ with respect $\bm{M}$.
+Definition 2.1 (Task Representation Learning).
+Given an input context variable $\bm{X}\in\mathcal{X}$ and its associated task/MDP random variable $\bm{M}\in\mathcal{M}$, task representation learning in COMRL aims to find a sufficient statistics $\bm{Z}$ of $\bm{X}$ with respect to $\bm{M}$.
 In pure statistical terms, Definition 2.1 implies that an ideal representation $\bm{Z}$ is a mapping of $\bm{X}$ that captures the mutual information $I(\bm{X};\bm{M})$.
 We therefore define the following dependency structures in terms of directed graphical models:
-Definition 2.2.
-The dependency graphs of COMRL is given by Fig.
+Definition 2.2 (Causal Decomposition).
+The dependency graphs of COMRL are given by Fig.
 2, where $\bm{X}_{b}$ and $\bm{X}_{t}$ are the behavior-related $(\bm{s},\bm{a})$-component and task-related $(\bm{s}^{\prime},r)$-component of the context $\bm{X}$, with $\bm{X}=(\bm{X}_{t},\bm{X}_{b})$.
@@
 The only exception is when tasks differ in initial state distribution $\rho_{0}$ or transition dynamics $T$, in which case the state variable $\bm{S}$ also depends on $\bm{M}$.
-We therefore define it as the behavior-related component, which should be minimally causally related (dashed lines) to $\bm{M}$ and $\bm{Z}$.
-Moreover, when $\bm{X}_{b}$ is given, $\bm{X}_{t}$ is fully characterized by the transition function $T:(\bm{s},\bm{a})\rightarrow\bm{s}^{\prime}$ and reward function $R:(\bm{s},\bm{a})\rightarrow r$ of $M$, which should be maximally causally related (solid lines) to $\bm{M}$ and $\bm{Z}$ and therefore be defined as the task-related component.
-Accordingly, we name $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ and $I(\bm{Z};\bm{X}_{b})$ the primary and lesser causality in our problem respectively.
-With this prior knowledge and assumption, we present the central theorem of this paper:
-Theorem 2.3.
+We therefore define it as the behavior-related component, which should be weakly causally related (dashed lines) to $\bm{M}$ and $\bm{Z}$.
+Moreover, when $\bm{X}_{b}$ is given, $\bm{X}_{t}$ is fully characterized by the transition function $T:(\bm{s},\bm{a})\rightarrow\bm{s}^{\prime}$ and reward function $R:(\bm{s},\bm{a})\rightarrow r$ of $M$, which should be strongly causally related (solid lines) to $\bm{M}$ and $\bm{Z}$ and therefore be defined as the task-related component.
+Mathematically, we find that rewriting $\bm{X}=(\bm{X}_{t},\bm{X}_{b})$ induces a causal decomposition of the mutual information $I(\bm{Z};\bm{X})$ by the chain rule [32]:
+$I(\bm{Z};\bm{X})=I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+I(\bm{Z};\bm{X}_{b}).$ (7)
+We thereby name $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ and $I(\bm{Z};\bm{X}_{b})$ the primary and lesser causality in our problem respectively.
+With the setup above, we present the central theorem of this paper:
+Theorem 2.3 (Central Theorem).
 Let $\equiv$ denote equality up to a constant, then
@@
 See Appendix B. ∎
-Theorem 2.3 induces several key observations.
-Firstly, the FOCAL and CORRO objectives operate as an upper bound and a lower bound of the relevant information $I(\bm{Z};\bm{M})$ respectively.
+Theorem 2.3 reveals several key observations.
+Firstly, the FOCAL and CORRO objectives operate as an upper and a lower bound of $I(\bm{Z};\bm{M})$ respectively.
 Since one would like to maximize $I(\bm{Z};\bm{M})$ according to Definition 2.1, CORRO, which maximizes the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$, can effectively optimize $I(\bm{Z};\bm{M})$ with theoretical assurance.
 However, FOCAL which maximizes the upper bound $I(\bm{Z};\bm{X}_{t},\bm{X}_{b})$ provides no guarantee for $I(\bm{Z};\bm{M})$.
-Mathematically since $I(\bm{Z};\bm{X})=I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+I(\bm{Z};\bm{X}_{b})$, maximizing the FOCAL objective may instead significantly elevates the lesser causality $I(\bm{Z};\bm{X}_{b})$, which is undesirable since it contains spurious correlation between the task representation $\bm{Z}$ and behavior policy $\pi_{\beta}$.
+By Eq.
+7, maximizing the FOCAL objective may instead significantly elevate the lesser causality $I(\bm{Z};\bm{X}_{b})$, which is undesirable since it contains spurious correlation between the task representation $\bm{Z}$ and behavior policy $\pi_{\beta}$.
 This explains why FOCAL is less robust to context shift compared to CORRO.
-Secondly, CSRO as the latest COMRL algorithm, inherently optimizes a linear combination of the FOCAL and CORRO objectives.
-In the $0\leq\lambda\leq 1$ regime, the CSRO objective becomes a convex interpolation of the upper bound $I(\bm{Z};\bm{X})$ and the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ of $I(\bm{Z};\bm{M})$, which in essence, enforces a trade-off between the causal ($\bm{Z}$ with $T\,\&\,\rho_{0}$) and spurious ($\bm{Z}$ with $\pi_{\beta}$) correlation contained in $I(\bm{Z};\bm{X}_{b})$.
+Secondly, CSRO as the latest COMRL algorithm among the three, inherently optimizes a linear combination of the FOCAL and CORRO objectives.
+In the $0\leq\lambda\leq 1$ regime, the CSRO objective becomes a convex interpolation of the upper bound $I(\bm{Z};\bm{X})$ and the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ of $I(\bm{Z};\bm{M})$, which in essence, enforces a trade-off between the causal ($\bm{Z}$ with $T$, $\rho_{0}$) and spurious ($\bm{Z}$ with $\pi_{\beta}$) correlation contained in $I(\bm{Z};\bm{X}_{b})$.
 This accounts for the improved performance of CSRO compared to FOCAL and CORRO.
@@
 […]
-Suppose the task label $M$ is given during training and a total of $n_{M}$ training tasks $\{M^{i}\}_{i=1}^{n_{M}}$ are drawn from the task distribution $p(M)$ for meta-training.
-Under this supervised scenario, by substituting Eqn 8 into 7, we have
+Suppose a total of $n_{M}$ training tasks $\{M^{i}\}_{i=1}^{n_{M}}$ are drawn from the task distribution $p(M)$ with the task label $M$ given for meta-training.
+Under this supervised scenario, by substituting Eq.
+9 into 8, we have
 […]
@@
 $=-\mathbb{E}_{\bm{x},\bm{z}\sim q_{\bm{\phi}}(\bm{z}|\bm{x})}\left[\sum_{j=1}^{n_{M}}\mathbbm{1}(M^{i}=M)\log p_{\bm{\theta}}(M^{i}|\bm{z})\right].$ (11)
-Note that $\mathcal{L}_{\textup{UNICORN-SUP}}$ is convex and operates as a finite-sample approximation of $-I(\bm{Z};\bm{M})$, for which we give the following theorem:
+Note that $\mathcal{L}_{\textup{UNICORN-SUP}}$ is convex and operates as a finite-sample approximation of $-I(\bm{Z};\bm{M})$, for which we derive the following bound:
 Theorem 2.4 (Concentration bound for supervised UNICORN).
@@
 The supervised UNICORN has the merit of directly estimating and optimizing the real objective $I(\bm{Z};\bm{M})$, which requires explicit knowledge of the task label $M^{i}$ and a substantial amount of task instances according to Theorem 2.4.
-To trade-off computation and performance, we choose to sample 20 training tasks for all RL environments in our experiments.
+For better trade-off of computation and performance, we choose to sample 20 training tasks for all RL environments in our experiments.
 Self-supervised UNICORN  In practice, offline RL datasets may often be collected with limited knowledge of the task specifications or labels.
@@
 where $0\leq\alpha\leq 1$ is a hyperparameter.
-Implementing each term in Eqn 12 allows ample decision choices, such as the contrastive losses in Eqs.
+Implementing each term in Eq.
+13 allows ample design choices, such as the contrastive losses in Eqs.
 3, 4 and 6 or autoregressive generation via Decision Transformer [35, 36] or RNN [34].
 For demonstration, in this paper we employ a contrastive objective $\mathcal{L}_{\textup{FOCAL}}$ as in Eq.
-3 for $I(\bm{Z};\bm{X})$ while approximate $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ by reconstruction.
-By the chain rule of mutual information [35]:
+3 for estimating $I(\bm{Z};\bm{X})$ while approximate $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ by reconstruction.
+By the chain rule:
 $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ $=I(\bm{X}_{t};\bm{Z},\bm{X}_{b})-I(\bm{X}_{t};\bm{X}_{b})$
@@
 which induces a generative objective $\mathcal{L}_{\textup{recon}}\coloneqq-I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ by reconstructing $\bm{X}_{t}$ with a decoder network $p_{\bm{\theta}}(\cdot|\bm{z},\bm{x_{b}})$ conditioning on $\bm{Z}$ and $\bm{X_{b}}$.
-As a result, the proposed unsupervised UNICORN objective can be rescaled as Eqn 15.
+As a result, the proposed unsupervised UNICORN objective can be rescaled as Eq.
+16:
+$\mathcal{L}_{\textup{UNICORN-SS}}\coloneqq\mathcal{L}_{\textup{recon}}+\frac{\alpha}{1-\alpha}\mathcal{L}_{\textup{FOCAL}}.$ (16)
 The influence of the hyper-parameter $\frac{\alpha}{1-\alpha}$ is shown in Appendix C.4.
-$\mathcal{L}_{\textup{UNICORN-SS}}\coloneqq\mathcal{L}_{\textup{recon}}+\frac{\alpha}{1-\alpha}\mathcal{L}_{\textup{FOCAL}}.$ (15)
-We summarize our learning procedure in Fig.
+We illustrate our learning procedure in Fig.
 3 with pseudo-code in Algorithms 1 and 2.
 A holistic comparison of our proposed algorithms with related contextual meta-RL methods is shown in Table 1.
-The extra KL divergence can be interpreted as the result of a variational approximation to an information bottleneck [36, 37] that constrains the mutual information between $\bm{Z}$ and $\bm{X}$, which we found unnecessary in our offline setting (see ablation in Table 6.) Behavior regularized actor critic [38] is employed to tackle the bootstrapping error [39] for downstream offline RL implementation.
+The extra KL divergence in methods like VariBAD and PEARL can be interpreted as the result of a variational approximation to an information bottleneck [37, 38] that constrains the mutual information between $\bm{Z}$ and $\bm{X}$, which we found unnecessary in our offline setting (see ablation in Table 6.) Behavior regularized actor critic [39] is employed to tackle the bootstrapping error [40] for downstream offline RL implementation.
 Table 1: Comparison between UNICORN instantiations and related existing contextual meta-RL methods.
@@
 3.1 Experimental Setup
-We adopt MuJoCo [25] and MetaWorld [3] benchmark to evaluate our method, involving six robotic locomotion tasks which require adaptation across reward functions (HalfCheetah-Dir, HalfCheetah-Vel, Ant-Dir, Reach) or across dynamics (Hopper-Param, Walker-Param).
-We compare UNICORN with six competitive OMRL algorithms.
-These baselines are categorized as context-based, gradient-based and transformer-based methods:
+Data Collection  Following FOCAL, for each task, we train a SAC [43] agent from scratch and collect its replay buffer as the offline dataset.
+Hence our default data represent a mixed distribution of behavior policies, ranging from random to expert level.
+Policy checkpoints are also kept for creating various new datasets in Sections 3.3 and 3.4.
+We adopt MuJoCo [26] and MetaWorld [3] benchmarks to evaluate our methods, which involve six robotic locomotion environments that require adaptation across reward functions (HalfCheetah-Dir, HalfCheetah-Vel, Ant-Dir, Reach) or across dynamics (Hopper-Param, Walker-Param).
+We compare UNICORN with six competitive OMRL algorithms, which are categorized as context-based, gradient-based and transformer-based methods:
 Table 2: Average testing returns of UNICORN against baselines on datasets collected by IID and OOD behavior policies.
-Each result is averaged by 6 random seeds.
+Results are averaged by 6 random seeds.
 The best is bolded and the second best is underlined.
@@
 Supervised is a context-based method, directly using actor-critic loss to train the policy/value networks and task encoder end-to-end. We find it a competitive baseline across all benchmarks.
-MACAW [12] is a gradient-based method, using supervised advantage-weighted regression for both the inner and outer loop of meta-training.
+MACAW [13] is a gradient-based method, using supervised advantage-weighted regression for both the inner and outer loops of meta-training.
 Prompt-DT [44] is a transformer-based method, taking context as the prompt of a Decision Transformer (DT) [35, 36] to solve OMRL as a conditional sequence modeling problem.
@@
 Solid curves refer to the mean performance of trials over 6 random seeds, and the shaded areas characterize the standard deviation of these trials.
-We perform our evaluation in a few-shot manner.
+Our evaluations are performed in a few-shot manner.
 For gradient-based methods, we first update the meta-policy with a batch of testing context and then evaluate the agent.
-For transformer-based methods, we utilize the trajectory segment as the prompt to condition the auto-regressive rollout.
+For transformer-based methods, we utilize trajectory segments as the prompt to condition the auto-regressive rollout.
 3.2 Few-Shot Generalization to In-Distribution Data
-Denote by $\{\pi_{\beta}^{i,t}\}_{i=1:N_{M},t=0:N}$ the checkpoints of behavior policies for all tasks logged during data collection (i.e., training SAC agents), where $i$ refers to the task index and $t$ represents training iteration.
-In the in-distribution (IID) setting, for each test task $M^{j}$, we use behavior policies only from the same tasks $\{\pi_{\beta}^{j,t}\}_{j,t=0:N}$ to collect the context and the conditioned rollouts, as visualized in Figure 1.
-Figure 4 illustrates the learning curves of UNICORN vs.
-all baselines, which correspond to the IID entries of Table 2.
+For each environment, the meta-policy is trained on 20 random tasks.
+To test IID generalization, we sample one trajectory for each testing task as the context, which enables few-shot adaptation of the agent via an updated task representation $\bm{z}$.
+Fig.
+4 illustrates the learning curves of UNICORN vs.
+all baselines in terms of testing returns, which correspond to the IID entries of Table 2.
 The results demonstrate that UNICORN variants, especially UNICORN-SS, consistently achieve SOTA performance across all benchmarks.
@@
 Although the gradient-based MACAW and the transformer-based Prompt-DT show faster convergence, their asymptotic performance leaves much to be desired.
-Moreover, the average training time for MACAW is almost three times longer than the context-based methods.
+Moreover, the average training time for MACAW is about three times longer than the context-based methods.
 3.3 Few-Shot Generalization to Out-of-Distribution Behavior Policies
-To evaluate OOD generalization, for each test task $M^{j}$, we use all behavior policies in $\{\pi_{\beta}^{i,t}\}_{{i=1:N_{M},t=0:N}}$ to collect the OOD contexts and the conditioned rollouts, as visualized in Figure 1.
-The OOD testing performance is measured by averaging returns across all testing tasks, as shown in the OOD entries of Table 2.
-While all methods suffer notable decline when facing OOD context, UNICORN maintains the strongest performance by a large margin.
+Denote by $\{\pi_{\beta}^{i,t}\}_{i=1:n_{M},t=0:n}$ the checkpoints of behavior policies for all tasks logged during data collection, where $i$ refers to the task index and $t$ represents training iteration.
+To evaluate OOD generalization, for each test task $M^{j}$, we use all behavior policies in $\{\pi_{\beta}^{i,t}\}_{{i=1:n_{M},t=0:n}}$ to collect the OOD contexts and the conditioned rollouts.
+The OOD testing performance is measured by averaging returns across all testing tasks and behavior policies, as shown in the OOD entries of Table 2.
+While all methods suffer notable decline when facing OOD context, UNICORN maintains the strongest performance by a significant margin.
 Another observation is that the context-based methods are in general significantly more robust compared to the gradient-based MACAW and transformer-based Prompt-DT.
-This also delineates the storyline of COMRL development from 2021 to date
+This also justifies the storyline of COMRL development from 2021 to date
 $\underset{(2021)}{\text{FOCAL}}\rightarrow\underset{(2022)}{\text{CORRO}}\rightarrow\underset{(2023)}{\text{CSRO}}\rightarrow\underset{(2024)}{\text{UNICORN}}$
@@
 Notably, despite CSRO having better IID results than UNICORN on Ant-Dir under the mixed distribution, UNICORN surpasses CSRO by a large margin under the narrow distribution (medium and expert) and random distribution.
-Table 3: UNICORN vs. baselines on Ant-Dir datasets of various qualities. Each result is averaged by 6 random seeds. The best is bolded and the second best is underlined.
+Table 3: UNICORN vs. baselines on Ant-Dir datasets of various qualities. Each result is averaged across 6 random seeds. The best is bolded and the second best is underlined.
 Algorithm Random Medium Expert IID OOD IID OOD IID OOD UNICORN-SS 81±18 62±6 220±23 243±10 279±10 262±13 UNICORN-SUP 75±15 60±5 140±11 126±32 247±15 229±19 CSRO 2±3 0±1 166±10 198±17 252±39 202±45 CORRO 1±1 0±0 8±5 -7±2 -4±10 -14±9 FOCAL 67±26 44±10 171±84 187±86 229±42 246±20 Supervised 65±6 47±12 149±50 110±80 249±33 215±60 MACAW 3±1 0±0 28±2 1±1 88±43 1±1 Prompt-DT 1±0 0±0 2±4 0±1 78±15 1±2
-We observe that CORRO fails on all three datasets.
-This may be due to the reliance of CORRO on the negative sample generator.
-Under narrow data distribution or poor random data distribution, it might be more challenging to generate negative samples of good quality.
+We observe that CORRO fails on all three datasets, which is likely attributable to the reliance of CORRO on the negative sample generator.
+On one hand, under narrow data distribution, there might be little overlap among high-density state-action supports of different tasks.
+On the other hand, samples collected by random policies are often indistinguishable across tasks [22].
+Both factors make it more challenging to train a robust generator for producing high quality negative samples conditioning on specific state-action pairs.
 As for Prompt-DT, its performance improves significantly with increased data quality, which is expected since the decision transformer adopts a behavior-cloning-like supervised learning style.
@@
 As UNICORN tackles task representation learning from a general information theoretic perspective, it is in principle model-agnostic.
-Hence a straightforward idea is to transfer UNICORN to other model architectures like DT, which we envision to be a promising backbone for large-scale training of RL foundation models.
-We employ a simple implementation by embedding the task representation vector $\bm{z}$ as the first token in sequence to prompt the learning and generation of DT.
-Instead of the unlearnable prompt used in the original Prompt-DT, we enforce task representation learning of the prompt by optimizing $\mathcal{L}_{\text{UNICORN-SS}}$, $\mathcal{L}_{\text{UNICORN-SUP}}$ and $\mathcal{L}_{\text{FOCAL}}$.
-We name these variants as UNICORN-SS-DT, UNICORN-SUP-DT and FOCAL-DT respectively.
-As shown in Table 4 and Figure 7, applying UNICORN and FOCAL on DT results in significant improvement in both IID and OOD generalization, with UNICORN being the superior option.
-Despite the gap between our implementation of UNICORN-DT and its MLP counterpart in terms of asymptotic performance, we expect UNICORN-DT to extrapolate favorably in the regime of greater dataset size and model parameters due to the scaling law of transformer [43] and DT [44].
+Hence a natural idea is to transfer UNICORN to other model architectures like DT, which we envision to be a promising backbone for large-scale training of RL foundation models.
+We employ a straightforward implementation by embedding the task representation vector $\bm{z}$ as the first token in sequence to prompt the learning and generation of DT, akin to in-context learning for large language models [45] and in-context RL [46, 47, 48, 49].
+Instead of the non-trainable prompt used in the original Prompt-DT, we enforce task representation learning of the prompt by optimizing $\mathcal{L}_{\text{UNICORN-SS}}$, $\mathcal{L}_{\text{UNICORN-SUP}}$ and $\mathcal{L}_{\text{FOCAL}}$.
+We name these variants UNICORN-SS-DT, UNICORN-SUP-DT and FOCAL-DT respectively.
+As shown in Table 4 and Fig.
+7, applying UNICORN and FOCAL on DT results in significant improvement in both IID and OOD generalization, with UNICORN being the superior option.
+Despite the gap between our implementation of UNICORN-DT and its MLP counterpart UNICORN-SS in terms of asymptotic performance, we expect UNICORN-DT to extrapolate favorably in the regime of greater dataset size and model parameters due to the scaling law of transformer [50] and DT [51].
 We leave this verification to future work.
 4.2 Can UNICORN be Exploited for Model-Based Paradigms?
-Figure 5: Testing returns for OOD tasks. The learning curve is averaged by 6 random seeds.
-With decoder $p_{\theta}(\bm{x}_{t}|\bm{z},\bm{x}_{b})$, UNICORN-SS can potentially enable a model-based paradigm by generating data using customized latent $\bm{z}$, which can be especially useful for generalization to OOD tasks.
-We implement this idea on Ant-Dir by sorting the tasks according to the polar angle of their goal direction ($\theta\in[0,2\pi)$), and construct task-level OOD by taking 20 consecutive tasks in the middle for training and the rest for testing (Figure 1).
+Figure 5: Testing returns for OOD tasks. The learning curves are averaged over 6 random seeds.
+With decoder $p_{\theta}(\bm{x}_{t}|\bm{z},\bm{x}_{b})$ as a world model [52], UNICORN-SS can potentially enable a model-based paradigm by generating data using customized latent $\bm{z}$, which can be especially useful for generalization to OOD tasks.
+We implement this idea on Ant-Dir by sorting the tasks according to the polar angle of their goal direction ($\theta\in[0,2\pi)$), and construct task-level OOD by taking 20 consecutive tasks in the middle for training and the rest for testing (Fig.
+1).
 Gaussian noise $\bm{\epsilon}$ is applied to the real task representation $\bm{z}$ and $(\bm{s},\bm{a},\bm{z}+\bm{\epsilon})$ is then fed to the decoder to generate $(\bm{s}^{\prime},r)$.
@@
 We adopt the conventional ensemble technique in model-based RL  [53, 54] to stabilize the training process.
-As shown in Figure 5, UNICORN-SS-enabled model-based RL is the only method to achieve positive performance in this extremely challenging scenario where context shift is induced by OOD tasks.
+As shown in Fig.
+5, UNICORN-SS-enabled model-based RL is the only method to achieve positive performance (indicating positive few-shot transfer) in this extremely challenging scenario where context shift is induced by OOD tasks.
 5 Conclusion & Limitation
-In this paper, we unify three major context-based offline meta-RL algorithms, FOCAL, CORRO and CSRO (potentially a lot more) into a single framework from an information theoretic representation learning perspective.
+In this paper, we unify three signature context-based offline meta-RL algorithms, FOCAL, CORRO and CSRO (potentially a lot more) into a single framework from an information theoretic representation learning perspective.
 Our theory offers valuable insight on how these methods are inherently connected and incrementally evolved in pursuit of better generalization against context shift.
 Based on the proposed framework, we instantiate two novel algorithms called UNICORN-SUP and UNICORN-SS which are demonstrated to be remarkably robust and broadly applicable through extensive experiments.
-We believe our study offers potential for new implementations, optimality bounds and algorithms for both fully-offline and offline-to-online COMRL paradigms.
-Mathematically, we expect our framework to be able to incorporate almost all existing COMRL methods that focus on task representation learning.
-Since according to Definition 2.1, as long as the method tries to solve COMRL by learning a sufficient statistic of w.r.t , it will eventually come down to optimizing an information-theoretic objective equivalent to $I(\bm{Z},\bm{M})$, or a lower/upper bound like the ones introduced by Theorem 2.3, up to some regularizations or constraints.
+We believe our study offers potential for new implementations, optimality bounds and algorithms for both fully-offline and offline-to-online COMRL.
+Moreover, we expect our framework to incorporate almost all existing COMRL methods that tackle task representation learning.
+Since according to Definition 2.1, as long as the method tries to solve COMRL by learning a sufficient statistic $\bm{Z}$ of $\bm{X}$ w.r.t.
+$\bm{M}$, it will eventually come down to optimizing an information-theoretic objective equivalent to $I(\bm{Z},\bm{M})$, or a lower/upper bound approximation like the ones introduced by  Theorem 2.3, up to some regularizations or constraints.
+With such mathematical generalization as well as the empirical evidence of UNICORN being model-agnostic (see Section 4.1), we envision it as a nascent offline pre-training paradigm of foundation models for decision making [21].
 Limitation  Since our framework assumes a decoupling of task representation learning and offline policy optimization, it does not directly elucidate how high-quality representations are indicative of higher downstream RL performance.
 Another limitation would be the scale of the experiments.
-Our offline datasets cover at most 40 tasks and $\sim 450k$ transitions, which might be the reason why UNICORN-SUP is inferior to UNICORN-SS ($n_{M}$ too small in Theorem 2.4).
+Our offline datasets cover at most 40 tasks and $\sim 450k$ transition tuples, which might be the reason why UNICORN-SUP is inferior to UNICORN-SS ($n_{M}$ too small to reduce the approximation error in Theorem 2.4).
 However, we expect our conclusion to extrapolate to larger task sets, dataset sizes and model parameters, for which the DT variants may demonstrate better scaling.
 Lastly, in principle, the information theoretic formalization in this work should be directly applicable to online RL.
-However, many of our key derivations rely on the static assumption of $\bm{M}$ and $\bm{X}$ (see Appendix B), which are apparently violated in the online scenario.
+However, many of our key derivations rely on the static assumption of $\bm{M}$ and $\bm{X}$ (see Appendix B for details), which are apparently violated in the online scenario.
 Extending our framework to online RL is interesting and nontrivial, which we leave for future work.
@@
 Remark: Consider a practical implementation of $p(\bm{z}|M)$ as a multivariate Gaussian $\mathcal{N}(\bm{\mu}_{M},\bm{\Sigma}_{M})$, which gives $H(\bm{Z}|\bm{M})=\frac{1}{2}\log[(2\pi e)^{\dim Z}|\bm{\Sigma}_{M}|]$ [32].
-Substituting into Eqn 20, we have with probability at least $1-\delta$,
+Substituting into Eqn 21, we have with probability at least $1-\delta$,
 $\left|\hat{I}(\bm{Z};\bm{M})-\bar{I}(\bm{Z};\bm{M})\right|\leq\sqrt{\frac{\textup{Var}(\log|\bm{\Sigma}_{M}|)}{2n_{M}\delta}}.$ (22)